### v2

# Costly Persuasion by a Partially Informed Sender

[…]
Costly Persuasion by a Partially Informed Sender
Shaofei Jiang The University of Bonn (email: sjiang@uni-bonn.de). I thank Sarah Auster, V. Bhaskar, Francesc Dilmé, William Fuchs, Sven Rady, Vasiliki Skreta, Maxwell Stinchcombe, Caroline Thomas, Mark Whitmeyer, Thomas Wiseman, and seminar participants at Bonn, UT Austin, SEA Annual Meeting, Midwest Economic Theory Conference, Stony Brook Game Theory Conference, AMES China, Asia Meeting of the Econometric Society, EEA-ESEM Congress, European Winter Meeting of the Econometric Society, SAET Conference for helpful comments. I gratefully acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - GZ 2047/1, Projekt-ID 390685813. Errors are my own.
(February 6, 2024)
[…]
I study a model of costly Bayesian persuasion by a privately and partially informed sender who conducts a public experiment. The cost of running an experiment is the expected reduction of a weighted log-likelihood ratio function of the sender’s belief. This is microfounded by a Wald sequential sampling problem where good news and bad news cost differently. I focus on equilibria that satisfy the D1 criterion. The equilibrium outcome depends crucially on the relative costs of drawing good and bad news in the experiment. If bad news is more costly, there exists a unique separating equilibrium, and the receiver unambiguously benefits from the sender’s private information. If good news is more costly, the single-crossing property fails. There may exist pooling and partial pooling equilibria, and in some equilibria, the receiver strictly suffers from sender private information.
Keywords: Bayesian persuasion, Informed principal, Signaling games
JEL Codes: C72, D82, D83
1. Introduction
Persuasion through public experimentation is prevalent. For example, a pharmaceutical company must preregister clinical trials and submit complete results to the Food and Drug Administration (FDA) as a part of the drug approval process. However, before conducting clinical trials, the pharmaceutical company already has more information about the drug than the FDA. Sources of the pharmaceutical company’s private information include internal research on related drugs, seeding trials and animal testings. [^1] Before entering clinical trials, a new drug undergoes preclinical (animal) testing, and the results are disclosed to the FDA in an Investigational New Drug Application (IND) for review. There is no mandate on disclosing complete preclinical results in INDs, and even in published preclinical research, unclear and non-transparent reporting practices (e.g., undisclosed statistical tests, undisclosed sample size) are prevalent (Gosselin,, 2021). A popular belief is that pharmaceutical companies’ private information puts the FDA at a disadvantage, and conversely, making this information public is socially beneficial as the FDA will be better informed. This has led to calls for more transparency throughout the pharmaceutical industry beyond just during clinical research.
Another example is startup funding. An entrepreneur develops a prototype in order to convince an investor that a new technology can be deployed in a marketable product. The prototype is developed according to plans submitted to the investor in advance, and the investor can independently evaluate whether the prototype meets set expectations. Despite the transparency in this process, the investor is naturally concerned that the entrepreneur has better knowledge about the technology and may capitalize on her private information when designing the prototype. Hence, investors often rely on external experts and carry out technical due diligence to mitigate risk.
However, it is not immediately clear whether these concerns are justified–that is, whether private information undermines the effectiveness of public experimentation and reduces welfare. On the one hand, if the pharmaceutical company designs different clinical trials based on its internal research, the FDA is able to infer the company’s private information from the design of the clinical trial and disentangle any effect of the private information. On the other hand, it is unclear whether in equilibrium, the pharmaceutical company chooses a more informative or less informative clinical trial compared to the counterfactual where it does not possess private information. Depending on the answers to these questions, private information may be beneficial or detrimental.
We study in this paper a costly persuasion game between a sender (e.g., the pharmaceutical company/the entrepreneur) and a receiver (e.g., the FDA/the investor). The sender persuades the receiver about an unknown, binary state of the world (e.g., whether a drug is effective/whether a technology is reliable) by conducting a public, costly experiment (e.g., a clinical trial/a prototype). The sender is privately and partially informed about the state of the world. At the outset of the game, she privately observes a noisy signal (e.g., results of internal research/proprietary knowledge about the technology), which is her type. Different types of the sender thus have different prior beliefs about the state and can choose different experiments. The receiver takes a binary action (e.g., whether to approve the drug/invest in the technology), and his optimal action depends on the state. Hence, this is a signaling game where the receiver tries to infer the state of the world from the sender’s choice of experiment as well as from the outcome of the experiment.
A key feature of the model is that the sender’s cost of running an experiment is type dependent. This feature is ubiquitous when using any posterior separable cost function to model experiments. In particular, we focus on the log-likelihood ratio cost, that is, the cost of running an experiment equals the expected reduction of a weighted log-likelihood ratio function evaluated at the sender’s belief. This cost function is microfounded via a Wald, (1945) sequential sampling problem, where the sender sequentially acquires public signals, at a cost, about the state of the world. The cost of drawing a signal is ex ante a random variable whose value depends on the signal’s realization. An experiment in the persuasion game is equivalent to a threshold stopping rule in the Wald sampling problem, and the cost of running an experiment is the expected cost of drawing signals before stopping.
Both the pharmaceutical example and the startup funding example can be modeled in this sequential manner, but with an important distinction. In the pharmaceutical example, a clinical trial consists of tests on individual patients, and the outcome of each patient is a signal about the effectiveness of the drug. From the point of view of persuading the FDA, it is good news if a patient recovers, and it is bad news if a patient does not recover. For the pharmaceutical company, bad news also leads to a higher cost of the clinical trial, since the company has to treat the patient using existing drugs if she does not recover. In the startup funding example, the process of developing a prototype can be viewed as developing a series of features. From the point of view of persuading the investor, successfully developing a feature is good news, and failure to do so is bad news. However, good news leads to a higher cost of developing the prototype because of moral hazard: to incentivize the engineers who work on the prototype, the entrepreneur needs to pay them a bonus atop their fixed wages if a feature is successfully developed.
Our main result is that the equilibrium outcome of the persuasion game depends on the relative costs of drawing good news and bad news in the experiment. As is common in signaling games, multiple equilibria exist, so we focus on equilibria that satisfy the D1 criterion (Banks and Sobel,, 1987). If bad news is more costly, there exists a unique separating equilibrium. Different sender types choose different experiments, hence the sender’s choice of experiment reveals her type. Moreover, sender private information benefits the receiver. That is, the receiver’s equilibrium payoff is higher than that in a benchmark model where the sender’s signal is public. This is because the sender chooses a more Blackwell informative experiment than she would have chosen in the benchmark model, in order to deter sender types whose prior beliefs are lower from mimicking. For example, a pharmaceutical company whose internal research shows favorable results will choose a larger scale clinical trial, which provides more information about the drug to the FDA. Hence, mandating disclosure of pharmaceutical companies’ internal research reduces public welfare, as it reduces pharmaceutical companies’ incentive to run informative clinical trials.
In contrast, if good news is more costly than bad news, running a more informative experiment costs more for a sender whose prior belief is higher, compared to a sender with a lower prior belief. Our model becomes a signaling game where the single-crossing property fails. Indeed, two indifference curves intersect up to three times. The equilibrium outcome is in general not unique, and the receiver may not learn the sender’s private information from her choice of experiment. In an equilibrium, an interval of sender types choose the same experiment (this includes pooling and separating equilibria as special cases), while other sender types each chooses a distinct experiment. We fully characterize the set of pooling equilibria and show that, in at least some pooling equilibria, the receiver’s payoff is strictly lower compared to that in a benchmark model where the sender observes no private signal. This, in turn, is bounded by the receiver’s payoff in the benchmark where the sender’s signal is public. Hence, entrepreneurs’ initial private information can hurt investors, and due diligence that helps reduce information asymmetry is necessary even when the process of developing prototypes is transparent.
The result that the single-crossing property may fail is not specific to the log-likelihood ratio costs we use. We show that the sender’s payoff satisfies double-crossing if experiments have the Shannon entropy cost.
1.1. Related literature
Kamenica and Gentzkow, (2011) introduce the study of Bayesian persuasion via unrestricted, costless experiments where the sender and the receiver have the same prior belief about the state of the world. Their main result is concavification. [^2] The sender’s payoff can be expressed as a value function over the (common) posterior belief, and the sender’s equilibrium payoff as a function of the prior is the concave closure of that value function. Alonso and Câmara, (2016) study an extension where the sender and the receiver have heterogeneous priors, but they agree to disagree. They derive a bijection between the sender’s and the receiver’s posterior beliefs, hence concavification can be applied after a translation of beliefs. Similarly, Gentzkow and Kamenica, (2014) relaxes the assumption that experiments are costless. With posterior separable (Caplin et al.,, 2022) costs of experiments, they show that a persuasion game with costly experiments and common priors is equivalent to one with costless experiments and heterogeneous priors. Hence, the equilibrium can be solved using concavification.
The study of persuasion games with a privately informed sender is more recent. Perez-Richet, (2014) studies equilibrium refinement in a persuasion game where the sender is fully informed of a binary state. Koessler and Skreta, (2023) study a more general information design problem by a fully informed designer, allowing for many agents and private messages. Alonso and Câmara, (2018) study persuasion by a partially informed sender who chooses an experiment from a restricted set of experiments. They provide conditions for a sender to never gain by becoming informed. All three papers assume that information transmission is costless and apply a generalized version of the inscrutability principle (Myerson,, 1983), which says that pooling equilibria are without loss of generality. In the current paper, in addition to introducing a cost of experiments, we assume that the public experiment cannot correlate with the sender’s private information conditional on the state of the world. Hence, the inscrutability principal does not apply. Two other papers that feature both costly experiments and sender private information are by Li and Li, (2013) and by Degan and Li, (2021). In both papers, a privately informed sender chooses from a restricted class of noisy signals that differ only on their precision, and the cost is increasing in the precision. In contrast, our paper allows the sender to choose any Blackwell experiment, and the cost of a Blackwell more informative experiment is higher.
Within the literature of Bayesian persuasion, the closest to the current paper is by Hedlund, (2017). He studies persuasion by a sender who is partially informed about a binary state and also focuses on equilibria satisfying the D1 criterion. However, experiments are assumed to be costless, and the sender’s payoff is strictly increasing in the receiver’s posterior belief. The equilibrium outcomes are either separating or fully disclosing. The introduction of costs of experiments adds rich dynamics to persuasion games. The sender’s type is payoff relevant, and the single-crossing property fails if good news is sufficiently more costly than bad news. As a result, pooling equilibria (which are not fully disclosing) and partial pooling equilibria may exist. In contrast, when experiments are costless, the single-crossing property is always satisfied.
The failure of single-crossing is worth noting beyond the Bayesian persuasion literature, since single-crossing is widely assumed in signaling games (see, e.g., the analysis of insurance markets by Rothschild and Stiglitz, (1976); Wilson, (1977)). It also plays an important role in applying various equilibrium refinements in signaling games (e.g., Riley, (1979); Cho and Kreps, (1987); Cho and Sobel, (1990)) and for monotone comparative statics (Milgrom and Shannon,, 1994). Chen et al., (2022) study signaling with a continuum of sender types where the sender’s preference exhibits double-crossing instead of single-crossing. They characterize equilibria satisfying the D1 criterion. In our model, when good news is more costly than bad news, two indifference curves of different sender types can intersect three times, hence violating even the double-crossing property.
Our paper is also related to the literature on the cost of information. Posterior-separable costs have been popular in modeling attention costs (e.g., Sims, (1998, 2003)) and are used to model costs of experiments by Gentzkow and Kamenica, (2014). However, an experiment, as is defined by Blackwell, (1953), is a concept independent of beliefs, and with heterogeneous priors, it is unclear which player’s beliefs should be used to compute the cost of an experiment. By studying a Wald, (1945) sequential sampling problem, we show that the cost of an experiment equals the expected reduction in a weighted log-likelihood ratio function evaluated at the sender’s belief. This is precisely the cost function studied by Pomatto et al., (2023), who provide an axiomatic foundation for log-likelihood ratio cost functions. Our setup of the sequential sampling problem is similar to that in Brocas and Carrillo, (2007) and Henry and Ottaviani, (2019), but we assume that the cost of acquiring each signal is a random variable. Our result complements other studies which microfound costs of information through sequential information acquisition (e.g., Morris and Strack, (2019); Bloedel and Zhong, (2020); Hébert and Woodford, (2023)).
The rest of the paper is organized as follows. Section 2 presents the game with a privately and partially informed sender and costly experiments. Section 3 studies a symmetric information benchmark where the sender and the receiver have heterogeneous priors. Section 4 presents the main theoretical results. We show that the single-crossing property fails when the cost of good news is sufficiently large compared to the cost of bad news, and we characterize the set of D1 pooling equilibrium outcomes when single-crossing fails. In Section 5, we characterize the unique separating equilibrium outcome when bad news is more costly. In Section 6, we present a Wald sampling problem which microfounds the log-likelihood ratio cost function. We also discuss partial pooling equilibria of the game and show that single-crossing also fails under the Shannon entropy cost. All proofs are in the Appendix.
2. The Model
There is a sender (she), and a receiver (he). At the outset of the game, Nature determines a binary state of the world $\omega\in\Omega:=\{G,B\}$ and a signal $\theta\in\Theta:=\{1,2,\dots,N\}$ according to a commonly known joint distribution $F$ with full support over $\Omega\times\Theta$. Let $\mu_{0}$ be the probability of the good state (i.e., $\omega=G$), and $\mu_{\theta}$ the probability of the good state conditional on the signal realization $\theta$. We assume that $10<\mu_{1}<\mu_{2}<\dots<\mu_{N}<$ 1. The sender privately observes the signal $\theta$, and neither player observes the state $\omega$. Therefore, $\theta$ is the sender’s type, and the sender’s prior belief on the good state is $\mu_{\theta}$. On the other hand, the receiver’s prior belief is $\mu_{0}$.
The game proceeds as follows. The sender publicly chooses an experiment $\pi$ on the state of the world. Without loss of generality, the experiment yields a binary outcome $s\in\{g,b\}$. That is, $\pi:\Omega\to\Delta(\{g,b\})$. The outcome of the chosen experiment $s$ is determined according to the distribution $\pi(\cdot|\omega)$ and is publicly observed. The receiver takes a binary action $a\in\{0,1\}$, and payoffs are realized. [^3] The model can be equivalently formulated as a mechanism design problem with an informed principal, where the outcome $g$ is an action recommendation for the receiver to take the action $a=1$, and the outcome $b$ is an action recommendation for the receiver to take the action $a=0$.
2.1. Strategies
Given an experiment $\pi$, let $p=\pi(g|G)$ and $q=\pi(g|B)$. Without loss of generality, $p\geq q$. An experiment is thus identified with the pair of probabilities $(p,q)$, and the set of feasible experiments is $\Pi=\{(p,q):1\geq p\geq q\geq 0\}$. We denote by $\Pi^{\circ}$ the interior of $\Pi$. An experiment $(p,q)$ is Blackwell more informative than another experiment $(p^{\prime},q^{\prime})$ if and only if $\frac{q}{p}\leq\frac{q^{\prime}}{p^{\prime}}$ and $\frac{1-p}{1-q}\leq\frac{1-p^{\prime}}{1-q^{\prime}}$.
A pure strategy of the sender $\{\pi_{\theta}\}_{\theta\in\Theta}$ is the collection of experiments chosen by all sender types, where $\pi_{\theta}\in\Pi$ is the experiment chosen by the type $\theta$ sender. A pure strategy of the receiver is $\mathbf{a}:\Pi\times\{g,b\}\to\{0,1\}$. It selects an action at every information set of the receiver, which is identified by the sender’s choice of experiment $\pi$ and its outcome $s$.
2.2. Beliefs
After observing the sender’s choice of experiment but before seeing its outcome, the receiver forms a belief about the sender’s type and the state of the world. Let $\gamma(\theta|\pi)$ denote his belief that the sender’s type is $\theta$ after experiment $\pi$ is chosen. Then $\beta(\pi):=\sum_{\theta\in\Theta}\gamma(\theta|\pi)\mu_{\theta}$ is the receiver’s interim belief on the good state. Notice that $\beta(\pi)\in[\mu_{1},\mu_{N}]$.
After the outcome is observed, both players update their beliefs. Let $\hat{\mu}(\theta,\pi,s)$ and $\hat{\beta}(\pi,s)$ be the posterior beliefs of the type $\theta$ sender and the receiver, respectively, that the state is good after observing outcome $s$ from experiment $\pi$.
2.3. Cost of experiments and the sender’s payoff
The sender strictly prefers the high receiver action over the low action. Her payoff $v(a,\pi|\theta)=a-c(\pi|\mu_{\theta})$ consists of two parts: a reward which is normalized to 1 if the receiver chooses the high action, minus the cost of the experiment $c(\pi|\mu_{\theta})$, which equals the expected reduction of a weighted log-likelihood ratio function evaluated at her belief. That is, the cost of running an experiment $\pi$ given the sender’s prior belief $\mu$ is
$c(\pi|\mu)=\mathbb{E}[H(\mu)-H(\hat{\mu})],$
where
$H(\mu)=C_{g}\mu\ln\left(\frac{1-\mu}{\mu}\right)+C_{b}(1-\mu)\ln\left(\frac{\mu}{1-\mu}\right),$
$C_{g},C_{b}>0$, and $\hat{\mu}$ is the sender’s posterior belief induced by the experiment $\pi$ [^4] The cost $c(\pi|\mu)\in\mathbb{R}_{+}\cup\{+\infty\}$ is well-defined for all $\mu\in(0,1)$ and $\pi\neq(0,0),(1,1)$. For completeness, let $c((0,0)|\mu)=c((1,1)|\mu)=0$ for all $\mu\in(0,1)$. Using the Kullback-Leibler divergence $D_{KL}$ (Kullback and Leibler,, 1951), the cost function can also be written as $c(\pi|\mu)=C_{g}\mu D_{KL}(P||Q)+C_{b}(1-\mu)D_{KL}(Q||P),$ where $P$ is the distribution of the experiment’s outcome in the good state (i.e., the Bernoulli distribution with success rate $p$), and $Q$ is that in the bad state (i.e., the Bernoulli distribution with success rate $q$). Intuitively, $D_{KL}(P||Q)$ and $D_{KL}(Q||P)$ measure the informativeness of the experiment $\pi$. Consider the hypothesis testing problem of $H_{0}:\omega=G$ versus $H_{1}:\omega=B$ given a large random sample of outcomes from the experiment $\pi$. A higher value of $D_{KL}(P||Q)$ implies a lower type II error probability, and a higher value of $D_{KL}(Q||P)$ implies a lower type I error probability (Cover and Thomas,, 2006).
A few remarks on the log-likelihood ratio cost function. Pomatto et al., (2023) show that this is the only family of cost functions satisfying three axioms. [^5] They consider cost functions as mappings from the the set of Blackwell experiments to the extended real line. The three axioms are: First, a Blackwell more informative experiment is more costly. Second, the cost of generating independent experiments is the sum of their individual costs. Third, the cost of generating an experiment with some probability is linear in the probability. We show in section 6.3 that the parameters $C_{g}$ and $C_{b}$ are the costs of drawing good and bad news, respectively. Hence, this cost function can accommodate both leading examples. The parameterization $C_{g}<C_{b}$ models scenarios such as pharmaceutical companies conducting clinical trials where bad news is more costly, while $C_{g}>C_{b}$ models scenarios such as entrepreneurs developing prototypes where good news is more costly. The cost of an experiment $c(\pi|\mu)$ is a linear function of $\mu$, but whether it is more costly for a more optimistic sender or a more pessimistic sender depends on the experiment. Finally, the cost of a Blackwell more informative experiment is always higher. The uninformative experiment [^6] An experiment is uninformative if $p=q$. Since all uninformative experiments are Blackwell equivalent and have zero cost, we will identify and refer to them as “the uninformative experiment.” has zero cost, and any experiment that reveals the state (i.e., $q<p=1$ or $0=q<p$) has an infinite cost.
A trade-off thus arises between information quality and cost for the sender. The sender wants to minimize the cost, but the experiment must be sufficiently informative so that the receiver is willing to take the high action if the outcome is good. In terms of $p$ and $q$, increasing $p$ improves information quality and increases the chance of getting the good outcome, but it also raises the cost. Increasing $q$ reduces the cost and increases the chance of getting the good outcome, but it reduces the informativeness of the experiment, causing the receiver to be less convinced after seeing the good outcome.
2.4. The receiver’s decision threshold
The receiver’s payoff $u(a,\omega)$ depends on his action $a$ and the state of the world $\omega$. By normalization, $u(0,G)=u(0,B)=0$, $u(1,G)=1$, and $u(1,B)=-\beta^{\star}/(1-\beta^{\star})$, hence $\beta^{\star}$ is the sequentially rational belief threshold of taking the high action. We assume that $\beta^{\star}\in(\mu_{N},1)$, so the receiver is never persuaded at the interim stage. The sender’s private signal is relatively noisy. Even if the receiver knows that the sender is the highest type, it is not optimal for him to take the high action without learning from the experiment.
We assume that the receiver follows a threshold decision rule and takes the high action if and only if his posterior belief is at least some threshold $\bar{\beta}\in(\beta^{\star},1)$. This threshold is common knowledge of the players and fixed throughout the paper. The assumption that $\bar{\beta}>\beta^{\star}$ comes from the receiver’s ability to commit to a higher standard in many relevant applications. The receiver is willing to make this commitment because, as is in the classic persuasion model, his equilibrium payoff is zero if he uses his sequentially rational belief threshold. By committing to a higher decision threshold, the receiver can achieve a positive payoff in equilibrium. This allows us to compare the receiver’s equilibrium payoffs with and without sender private information in a meaningful way.
2.5. Equilibrium
An equilibrium consists of pure strategies of the players, $\{\pi_{\theta}\}_{\theta\in\Theta}$ and $\mathbf{a}:\Pi\times\{g,b\}\to\{0,1\}$, and the receiver’s system of beliefs $\beta:\Pi\to[\mu_{1},\mu_{N}]$ and $\hat{\beta}:\Pi\times\{g,b\}\to[0,1]$, such that
  1. (1)
    Given the receiver’s strategy $\mathbf{a}$, the sender’s strategy is optimal, i.e.,
    $\pi_{\theta}\in\operatorname*{argmax}_{\pi\in\Pi}\mathbb{E}[v(\mathbf{a}(\pi,s),\pi|\theta)]$
    for all $\theta\in\Theta$;
  2. (2)
    The receiver follows the threshold rule, i.e., $\mathbf{a}(\pi,s)=1$ if and only if $\hat{\beta}(\pi,s)\geq\bar{\beta}$;
  3. (3)
    Beliefs are updated using Bayes’ rule whenever possible. That is,
    $\beta(\pi)=\frac{F(\{G\}\times\Theta_{\pi})}{F(\Omega\times\Theta_{\pi})}$
    if $\Theta_{\pi}:=\{\theta:\pi_{\theta}=\pi\}$ is nonempty, and
    $\hat{\beta}(\pi,s)=\mathbf{B}(\beta(\pi),\pi,s):=\frac{\beta(\pi)\pi(s|G)}{\beta(\pi)\pi(s|G)+(1-\beta(\pi))\pi(s|B)}$
    if $\pi(s|G)+\pi(s|B)\neq 0$.
We say an equilibrium is a persuasion equilibrium if the high action is taken with positive probability on the equilibrium path. Otherwise, it is a trivial equilibrium. A persuasion equilibrium is a pooling equilibrium if all sender types choose the same experiment. It is a separating equilibrium if, except for those who choose the uninformative experiment, all sender types choose distinct experiments. That is, $\pi_{i}\neq\pi_{j}$ for all $i\neq j$ and $\pi_{i},\pi_{j}\in\Pi^{\circ}$. Any persuasion equilibrium that is neither pooling nor separating is a partial pooling equilibrium. Given an equilibrium, we call the collection of experiments $\{\pi_{\theta}\}_{\theta\in\Theta}$ the equilibrium outcome. Two equilibrium outcomes $\{\pi_{\theta}\}_{\theta\in\Theta}$ and $\{\pi_{\theta}^{\prime}\}_{\theta\in\Theta}$ are equivalent if $\pi_{\theta}=\pi_{\theta}^{\prime}$ for all $\theta$ such that $\pi_{\theta}\in\Pi^{\circ}$ or $\pi_{\theta}^{\prime}\in\Pi^{\circ}$. The equilibrium outcome is unique if all equilibrium outcomes are equivalent.
2.6. The D1 criterion
We focus on equilibria that satisfy the D1 criterion (Banks and Sobel,, 1987). Given the receiver’s interim belief $\beta\in[\mu_{1},\mu_{N}]$, let $\bar{v}(\beta,\pi|\theta)$ be the type $\theta$ sender’s expected payoff from choosing an experiment $\pi$. That is,
$\bar{v}(\beta,\pi|\theta)=\mathbb{P}[\mathbf{B}(\beta,\pi,s)\geq\bar{\beta}]-c(\pi|\mu_{\theta}).$
Fixing an equilibrium, denote by $v_{\theta}^{\star}=\bar{v}(\beta(\pi_{\theta}),\pi_{\theta}|\theta)$ the sender’s equilibrium payoff, and for any deviation $\pi\in\Pi\setminus\{\pi_{\theta}\}_{\theta\in\Theta}$, let
$D_{\theta}(\pi)=\{\beta\in[\mu_{1},\mu_{N}]:\bar{v}(\beta,\pi|\theta)>v_{\theta}^{\star}\},$
$D^{0}_{\theta}(\pi)=\{\beta\in[\mu_{1},\mu_{N}]:\bar{v}(\beta,\pi|\theta)\geq v_{\theta}^{\star}\}.$
$D_{\theta}(\pi)$ is the set of receiver’s interim beliefs given which $\pi$ is a profitable deviation for the type $\theta$ sender, and $D^{0}_{\theta}(\pi)$ is the set of receiver interim beliefs given which the deviation gives the sender at least the same payoff as her equilibrium payoff. An important special case is when $D^{0}_{\theta}(\pi)$ is empty, that is, the deviation $\pi$ is strictly equilibrium dominated for the sender regardless of the receiver’s interim belief $\beta(\pi)$.
An equilibrium satisfies the D1 criterion if: for all pairs of distinct sender types $(i,j)$ and deviations $\pi$, if $D_{i}^{0}(\pi)\subsetneq D_{j}(\pi)$, then the receiver’s off-path interim belief $\beta(\pi)$ lies in the convex hull of $\{\mu_{\theta}:\theta\neq i\}$. In words, if some sender type $j$ is keener to deviate to $\pi$ than sender type $i$, in the sense that this deviation is profitable for her given a larger set of receiver beliefs than for sender type $i$, then the receiver should not attribute this deviation to sender type $i$.
3. Symmetric Information Benchmark
We first study how the sender’s trade-off between information quality and cost depends on her prior in the absence of the receiver’s interim learning. Consider a symmetric information game where the sender and the receiver have heterogeneous priors, but they agree to disagree. The sender publicly chooses an experiment, and the receiver takes an action based on the outcome of that experiment. The main result (Proposition 3.1) is that the informativeness of the equilibrium outcome is ranked by the sender’s prior belief, but whether a more optimistic sender chooses a more or less informative experiment depends on the relative costs of good news and bad news.
Let $\mu_{r}<\bar{\beta}$ be the receiver’s prior belief, and $\mu_{s}$ the sender’s prior belief. We say an experiment $\pi=(p,q)$ is persuasive at belief $\mu_{r}$ if the receiver will take the high action upon seeing the good outcome, that is, if $\mathbf{B}(\mu_{r},\pi,g)\geq\bar{\beta}$. Equivalently,
$\frac{q}{p}\leq\mathbf{Q}(\mu_{r}):=\frac{\mu_{r}}{1-\mu_{r}}\Big{/}\frac{\bar{\beta}}{1-\bar{\beta}}.$
All other experiments are unpersuasive at belief $\mu_{r}$. Following an unpersuasive experiment, the receiver always takes the low action.
The sender’s expected payoff from choosing an experiment $\pi$ that is persuasive at belief $\mu_{r}$ is
$f(\pi,\mu_{s}):=\mu_{s}p+(1-\mu_{s})q-c(\pi|\mu_{s}).$
It is strictly increasing in $q$. As $q$ increases, the probability that the receiver takes the high action increases, and the cost of the experiment decreases as it becomes less informative. Therefore, it is without loss to focus on experiments such that $q=\mathbf{Q}(\mu_{r})p$. On the other hand, the sender’s expected payoff from choosing an unpersuasive experiment $\pi$ is $-c(\pi|\mu_{r})\leq 0$, and the sender receives zero payoff if and only if the uninformative experiment is chosen.
Hence, the sender’s equilibrium payoff in the symmetric information benchmark is [^7] It is easy to verify that the objective function in (3.1) is right-continuous at $p=0$. Hence, the maximum exists and is nonnegative.
(3.1) $V(\mu_{s},\mu_{r}):=\max_{p\in[0,1]}f((p,\mathbf{Q}(\mu_{r})p),\mu_{s})\geq 0.$
A unique equilibrium exists, and $V(\mu_{s},\mu_{r})>0$ if and only if the cost of experiments is low in the sense that a weighted average of $C_{g}$ and $C_{b}$ is below a threshold (see Lemma A.1 for the precise statement). If the cost of experiments is low, the sender chooses an experiment that is persuasive at belief $\mu_{r}$. If the cost of experiments is high, persuasion is not possible in equilibrium. The sender chooses the uninformative experiment, and her equilibrium payoff is zero.
We now present the main comparative statics results for the symmetric information benchmark. Fixing the receiver’s prior belief, the informativeness of the equilibrium outcome is ranked by the sender’s prior belief. When bad news is more costly than good news, the sender chooses a Blackwell more informative experiment if her prior is higher. When good news is sufficiently more costly than bad news, the sender chooses a Blackwell less informative experiment if her prior is higher.
Proposition 3.1.
There exists $\hat{K}:\mathbb{R}_{++}\to\mathbb{R}_{++}$ such that the unique equilibrium outcome $\hat{\pi}$ of the symmetric information benchmark is independent of the sender’s prior $\mu_{s}$ if and only if $C_{g}=\hat{K}(C_{b})$. If $C_{g}>\hat{K}(C_{b})$, $\hat{\pi}$ is less Blackwell informative as $\mu_{s}$ increases; if $C_{g}<\hat{K}(C_{b})$, $\hat{\pi}$ is more Blackwell informative as $\mu_{s}$ increases. Moreover, $\hat{K}$ is twice differentiable, increasing and strictly concave. For all $C_{b}>0$, $\hat{K}(C_{b})>C_{b}$, and $\lim_{C_{b}\downarrow 0}\hat{K}^{\prime}(C_{b})=\infty$.
Figure 3.1 illustrates the results in the space of cost parameters $(C_{b},C_{g})$. Given the players’ prior beliefs $\mu_{r}$ and $\mu_{s}$, persuasion is possible if and only if the cost parameters are below the solid, downward sloping line. Fixing $\mu_{r}$ and increasing $\mu_{s}$, this line rotates around a fixed point, and its slope decreases in absolute value. The highlighted curve plots the function $\hat{K}$, which is independent of $\mu_{s}$ or $\mu_{r}$. If $C_{g}=\hat{K}(C_{b})$, the unique equilibrium outcome is independent of the sender’s prior. If $C_{g}<\hat{K}(C_{b})$, persuasion is possible in equilibrium if the sender’s prior is sufficiently high, and the equilibrium outcome is more Blackwell informative as the sender’s prior increases. If $C_{g}>\hat{K}(C_{b})$, persuasion is possible in equilibrium if the sender’s prior is sufficiently low, and the equilibrium outcome is less Blackwell informative as the sender’s prior increases.
[…]
Figure 3.1: Cost parameters where persuasion is possible and comparative statics in the symmetric information benchmark
A special case where $\mu_{s}=\mu_{r}$ is studied by Gentzkow and Kamenica, (2014) and pertains to the welfare comparison in our model. We consider two benchmarks. In the no-signal benchmark, the sender does not observe the noisy signal $\theta$, hence the players have common prior $\mu_{0}$. In the public-signal benchmark, the noisy signal $\theta$ is observed by both the sender and the receiver. Hence, after observing the noisy signal, the continuation game is one with common prior $\mu_{\theta}$. Providing public information weakly benefits the receiver, and increasing the common prior weakly benefits the sender.
Lemma 3.2.
Let $U^{ns}$ be the receiver’s ex ante expected payoff in the no-signal benchmark, and $U^{ps}$ his payoff in the public-signal benchmark. $U^{ps}\geq U^{ns}$. Moreover, the sender’s equilibrium payoff in a symmetric information benchmark with common prior $\mu$, $V(\mu,\mu)$, is nondecreasing in $\mu$.
4. The Single-Crossing Property and Pooling Equilibria
We now turn to the model with sender private information. We define the sender’s marginal rate of substitution. The single-crossing property is satisfied if the marginal rate of substitution is weakly decreasing in the sender’s prior belief. In our model, single-crossing fails when good news is sufficiently more costly than bad news (Proposition 4.1). When the single-crossing property is satisfied, all D1 persuasion equilibria are separating. When single-crossing fails, pooling equilibria may be selected by D1, and Proposition 4.3 characterizes all pooling equilibrium outcomes.
4.1. Marginal rate of substitution and single-crossing
The marginal rate of substitution (of $p$ for $q$ for the sender type $\theta$) is
$MRS(\pi|\mu_{\theta})=-\frac{\partial f(\pi,\mu_{\theta})/\partial p}{\partial f(\pi,\mu_{\theta})/\partial q}.$
Supposing that the receiver takes the high action when he sees the good outcome, that is, the sender’s expected payoff is $f(\pi,\mu_{\theta})$, the marginal rate of substitution is the marginal benefit of increasing $p$ relative to increasing $q$. Recall that increasing either $p$ or $q$ increases the probability that the receiver takes the high action, but increasing $p$ incurs a higher cost, whereas increasing $q$ reduces the cost. Hence, the marginal rate of substitution is negative when $p$ is small, but it becomes positive and goes to infinity as $p$ goes to $1$.
The marginal rate of substitution has the usual geometric representation. In Figure 4.1, the right triangle is the set of experiments $\Pi$, where the conditional probabilities $p$ and $q$ are shown on the horizontal and vertical axes, respectively. The solid curve shows an indifference curve of some sender type $h$. The marginal rate of substitution $MRS(\pi|\mu_{h})$ is the slope of the indifference curve at the experiment $\pi$.
[…]
Figure 4.1: The single-crossing property and an example of a nearby deviation
At the uninformative experiment, the marginal rate of substitution
(4.1) $MRS(\pi|\mu_{\theta})=-\frac{\mu_{\theta}}{1-\mu_{\theta}}$
is decreasing in $\mu_{\theta}$. We say that the single-crossing property is satisfied if the marginal rate of substitution is weakly decreasing in the sender’s prior belief at all experiments, i.e.,
$\frac{\partial}{\partial\mu}MRS(\pi|\mu)\leq 0$
for all $\pi\in\Pi^{\circ}$. [^8] The marginal rate of substitution is not well defined if $p=1$ or $q=0$, hence the need to define single-crossing only on $\Pi^{\circ}$. Geometrically, this implies that the indifference curve of a lower sender type is everywhere steeper than that of a higher sender type, hence any two indifference curves intersect at most once. Figure 4.1 shows an example. Consider a lower sender type $l<h$, whose indifference curve through the experiment $\pi$ is the dashed curve. The indifference curve of the low-type sender $l$ has a more upward slope than that of the high-type sender $h$ at the experiment $\pi$, and the two indifference curves intersect only at $\pi$.
4.2. Failure of single-crossing
If experiments are costless, the single-crossing property is always satisfied, since the sender’s expected payoff $f(\pi,\mu_{\theta})=\mu_{\theta}p+(1-\mu_{\theta})q$ is simply the probability of the good outcome, and the marginal rate of substitution, given by (4.1), is decreasing in $\mu_{\theta}$. When experiments are costly, the sender’s trade-off is more involved, as the sender’s choice of experiment affects her payoff also through its cost. Proposition 4.1 below shows that the single-crossing property fails if good news is sufficiently more costly than bad news.
Proposition 4.1.
The single-crossing property is satisfied if and only if $C_{g}\leq\hat{K}(C_{b})$. If $C_{g}>\hat{K}(C_{b})$, there exist $\hat{\mathbf{p}},\check{\mathbf{p}}:(0,1)\to(0,1)$ such that $q<\hat{\mathbf{p}}(q)<\check{\mathbf{p}}(q)$ for all $q\in(0,1)$, and
(4.2) $\frac{\partial}{\partial\mu}MRS(\pi|\mu)\left\{\begin{array}[]{ll}<0&\text{if $p<\hat{\mathbf{p}}(q)$ or $p>\check{\mathbf{p}}(q)$}\\ =0&\text{if $p=\hat{\mathbf{p}}(q)$ or $p=\check{\mathbf{p}}(q)$}\\ >0&\text{if $\hat{\mathbf{p}}(q)<p<\check{\mathbf{p}}(q)$}\end{array}\right..$
Moreover, $\hat{\mathbf{p}}(q)$ is decreasing in $C_{g}$ and increasing in $C_{b}$, and $\check{\mathbf{p}}(q)$ is increasing in $C_{g}$ and decreasing in $C_{b}$ for all $q\in(0,1)$.
[…]
Figure 4.2: Experiments that are robust to nearby deviations when the single-crossing property fails
Figure 4.2 illustrates the results. When single-crossing fails, there exist loci of experiments at which the marginal rate of substitution is constant across sender types. These experiments are highlighted by the curves $p=\hat{\mathbf{p}}(q)$ and $p=\check{\mathbf{p}}(q)$. If an experiment is sufficiently uninformative (i.e., if $p<\hat{\mathbf{p}}(q)$) or sufficiently informative (i.e., if $p>\check{\mathbf{p}}(q)$), the marginal rate of substitution is decreasing in the sender’s type. In the interim region (i.e., if $\hat{\mathbf{p}}(q)<p<\check{\mathbf{p}}(q)$), the marginal rate of substitution is increasing in the sender’s type. Hence, two indifference curves can intersect up to three times, one in each region.
Consider two experiments $\pi_{1}=(p_{1},q_{1})$ and $\pi_{2}=(p_{2},q_{2})$ such that $p_{1}=\hat{\mathbf{p}}(q_{1})$ and $p_{2}=\check{\mathbf{p}}(q_{2})$. Figure 4.2 shows the indifference curves of two types of the sender $h>l$ at these experiments. At both experiments, the indifference curves are tangent. However, at experiment $\pi_{1}$, the high-type sender’s indifference curve is more convex than the low-type sender’s indifference curve. Hence, in a neighborhood of $\pi_{1}$, the high-type sender’s indifference curve is higher than the low-type sender’s indifference curve. But at experiment $\pi_{2}$, the low-type sender’s indifference curve is more convex and therefore higher than that of the high-type sender.
4.3. Nearby deviations
The marginal rate of substitution calibrates the sender’s incentive to make slight deviations. If two sender types have different marginal rates of substitution at an experiment, the D1 criterion can rule out any equilibrium where they both choose this experiment.
Lemma 4.2.
Let $\pi\in\Pi^{\circ}$. If $\frac{\partial}{\partial\mu}MRS(\pi|\mu)\neq 0$, any equilibrium such that $\pi_{i}=\pi_{j}=\pi$ for some $i\neq j$ does not satisfy the D1 criterion.
Let us illustrate this using Figure 4.1. For simplicity, assume that there are only two types of the sender $h$ and $l$, and they both choose $\pi$ in an equilibrium. The experiment $\pi$ must be persuasive at $\mu_{0}$, which is the receiver’s interim belief in the pooling equilibrium. Hence, we can select a nearby deviation $\tilde{\pi}$ that is persuasive at belief $\mu_{h}$. For the high-type sender, deviating to $\tilde{\pi}$ is profitable if the receiver’s interim belief $\beta(\tilde{\pi})$ is sufficiently high. Specifically, if the receiver believes that this deviation is made by the high-type sender, the sender gets $f(\tilde{\pi},\mu_{h})$, which is higher than her equilibrium payoff $f(\pi,\mu_{h})$. In contrast, this deviation is strictly equilibrium dominated for the low-type sender. [^9] Hence, the intuitive criterion (Cho and Kreps,, 1987) can also rule out the pooling equilibrium. It is generally true in our model that the intuitive criterion and the D1 criterion select the same equilibrium outcomes. Even if the receiver believes that this deviation is by the high-type sender, the low-type sender’s deviation payoff $f(\tilde{\pi},\mu_{l})$ is strictly less than her equilibrium payoff $f(\pi,\mu_{l})$. Therefore, the D1 criterion asserts that it is only reasonable to attribute the deviation to the high type, but this makes $\tilde{\pi}$ a profitable deviation for the high type. Therefore, there is no D1 pooling equilibrium where both types choose $\pi$.
As a result of Lemma 4.2, when the single-crossing property is satisfied, any D1 persuasion equilibrium is a separating equilibrium. When the single-crossing property fails, pooling or partial pooling equilibria may be selected by the D1 criterion. We say an experiment $\pi=(p,q)$ is robust to nearby deviations if $p=\hat{\mathbf{p}}(q)$. For example, at experiment $\pi_{1}$ in Figure 4.2, the high type sender’s indifference curve is lower than the low type sender’s indifference curve, hence any nearby deviation that is profitable for the high-type sender is also profitable for the low-type sender. Therefore, the D1 criterion is moot in determining the receiver’s beliefs after seeing these nearby deviations. In contrast, if $p=\check{\mathbf{p}}(q)$ (e.g., at experiment $\pi_{2}$), the high-type sender’s indifference curve is lower. Therefore, we can find a nearby deviation (e.g., $\tilde{\pi}$) that is profitable only for the high-type sender but strictly equilibrium dominated for the low-type sender. Therefore, $\pi_{2}$ cannot be chosen by more than one sender type in any D1 equilibrium.
4.4. Large deviations and pooling equilibria
A pooling equilibrium outcome that is robust to nearby deviations does not necessarily satisfy the D1 criterion. We are to illustrate in Figure 4.3 that there may be a large deviation that is profitable only for the high-type sender.
[…]
Figure 4.3: An example of a large deviation that breaks a pooling equilibrium
Let $\pi=(\hat{\mathbf{p}}(q),q)$ be a pooling equilibrium outcome. The indifference curves of the two sender types are tangent at $\pi$, but they intersect again to the right of $\pi$, at $\pi^{\prime}$, and the low-type sender’s indifference curve is steeper at $\pi^{\prime}$. Therefore, we can find a deviation close to $\pi^{\prime}$, say $\tilde{\pi}$, that is above the high-type sender’s indifference curve and below the low-type sender’s indifference curve. Two cases are possible. If $\tilde{\pi}$ is persuasive at belief $\mu_{h}$, then this deviation is profitable for the high-type sender for some interim belief of the receiver, but it is strictly equilibrium dominated for the low-type sender. In this case, the D1 criterion requires that the receiver attributes this deviation to the high sender type, thus breaking the pooling equilibrium. However, if $\tilde{\pi}$ is unpersuasive even at belief $\mu_{h}$–that is, the receiver always chooses the low action–then $\tilde{\pi}$ is not a profitable deviation for either type of the sender. In this case, the D1 criterion does not impose any restriction on the receiver’s interim belief and cannot rule out the pooling equilibrium outcome $\pi$.
We say an equilibrium $\pi=(\hat{\mathbf{p}}(q),q)$ is robust to large deviations if, for all experiments $\tilde{\pi}$ that are persuasive at belief $\mu_{N}$, $f(\tilde{\pi},\mu_{N})>f(\pi,\mu_{N})$ implies $f(\tilde{\pi},\mu_{\theta})>f(\pi,\mu_{\theta})$ for all $\theta\in\Theta$. That is, if a deviation is profitable for the highest sender type, it is profitable for all lower sender types, so the highest sender type cannot convince the receiver of her type by choosing an off the equilibrium path experiment.
A pooling equilibrium outcome satisfies the D1 criterion if and only if it is robust both to nearby deviations and to large deviations. The following proposition characterizes the set of all D1 pooling equilibrium outcomes.
Proposition 4.3.
Let $C_{g}>\hat{K}(C_{b})$. There exists $\hat{q}\geq 0$ such that $\pi=(p,q)$ is a D1 pooling equilibrium outcome if and only if: (i) $\frac{q}{p}\leq\mathbf{Q}(\mu_{0})$; (ii) $f(\pi,\mu_{1})\geq V(\mu_{1},\mu_{1})$; (iii) $p=\hat{\mathbf{p}}(q)$; (iv) $q\geq\hat{q}$.
Condition (i) is the receiver’s obedience constraint. Given his interim belief $\beta(\pi)=\mu_{0}$, the receiver must be willing to take the high action if $\pi$ reveals the good outcome. Condition (ii) is the participation constraint of the lowest sender type. Since $\mu_{1}$ is the lowest possible interim belief of the receiver, the lowest sender type can achieve at least the payoff in the symmetric information benchmark where the receiver’s prior is $\mu_{1}$, that is, $V(\mu_{1},\mu_{1})$, in every equilibrium. Conditions (iii) and (iv) ensure that the equilibrium outcome is robust to nearby deviations and large deviations, respectively. Notice that restricting to experiments such that $p=\hat{\mathbf{p}}(q)$, the participation constraint of the lowest sender type implies those of all sender types. This simplifies the process of solving for D1 pooling equilibria.
Proposition 4.3 has a few implications. First, when $\mu_{0}$ is small, $\frac{q}{\hat{\mathbf{p}}(q)}>\mathbf{Q}(\mu_{0})$ for all $q\in(0,1)$, so no D1 pooling equilibrium exists. Indeed, when $\mu_{0}$ is sufficiently small and $C_{g}\neq\hat{K}(C_{b})$, the game has no pooling equilibrium. Second, if a D1 pooling equilibrium exists, the set of D1 pooling equilibrium outcomes is a continuum. As both $\frac{q}{\hat{\mathbf{p}}(q)}$ and $f((\hat{\mathbf{p}}(q),q),\mu_{1})$ are increasing in $q$, conditions (i) is an upper bound on $q$, and (ii) is a lower bound. Together with condition (iv), an experiment $\pi=(\hat{\mathbf{p}}(q),q)$ is a D1 pooling equilibrium outcome if and only if $q\in[\underline{q},\bar{q}]$ for some $\underline{q}$ and $\bar{q}$ such that $\frac{\hat{\mathbf{p}}(\bar{q})}{\bar{q}}=\mathbf{Q}(\mu_{0})$. [^10] If $V(\mu_{1},\mu_{1})=0$, (ii) is satisfied by all $q>0$. If, in addition, $\hat{q}=0$, the set of D1 pooling equilibrium outcomes is $\{(\hat{\mathbf{p}}(q),q):q\in(0,\bar{q}]\}$. We discuss this case in further detail in Section 4.5. Finally, we observe that the set of D1 pooling equilibrium outcomes is not Blackwell ordered, and the receiver’s equilibrium payoff may not be monotonic in $q$. The following proposition shows that in at least some D1 pooling equilibria, the receiver’s payoff is strictly less than her payoff in the no-signal benchmark.
Proposition 4.4.
Suppose that the set of D1 pooling equilibrium outcomes $\{\pi=(\hat{\mathbf{p}}(q),q):q\in[\underline{q},\bar{q}]\}$ is nonempty, and denote by $U^{p}(q)$ the receiver’s payoff in an equilibrium where the sender chooses $(\hat{\mathbf{p}}(q),q)$. $U^{p}$ is concave, and $U^{p}(\bar{q})<U^{ns}$.
4.5. Trivial equilibria
There exists an equilibrium where all sender types choose the uninformative experiment if the cost of experiments is sufficiently high. Specifically, if $V(\mu_{1},\mu_{1})=0$, no sender type can achieve a positive payoff in the symmetric information benchmark with the receiver’s prior $\mu_{1}$. Therefore, there is a trivial equilibrium where the receiver attributes any deviation to the lowest sender type. A trivial equilibrium, however, may be ruled out by the D1 criterion using large deviations. The analysis is akin to that of pooling equilibria. There exists an experiment $\pi^{\prime}\in\Pi^{\circ}$ that gives all sender types zero payoffs, i.e., $f(\pi^{\prime},\mu_{\theta})=0$ for all $\theta$, and at $\pi^{\prime}$, the marginal rate of substitution is decreasing in the sender’s type. Hence, we can find a deviation $\tilde{\pi}$ such that $f(\tilde{\pi},\mu_{N})>0$, and $f(\tilde{\pi},\mu_{\theta})<0$ for all $\theta<N$. If $\tilde{\pi}$ is persuasive at belief $\mu_{N}$, the trivial equilibrium is ruled out by the D1 criterion. The following proposition shows that there exists a trivial D1 equilibrium if and only if $\mu_{N}$ is sufficiently small.
Proposition 4.5.
Let $C_{g}\geq\hat{K}(C_{b})$. There exists $\bar{\mu}\in(0,\bar{\beta})$ such that a trivial D1 equilibrium exists if and only if (i) $V(\mu_{1},\mu_{1})=0$, and (ii) $\mu_{N}\leq\bar{\mu}$. If $V(\mu_{N},\mu_{N})=0$, the unique D1 equilibrium outcome is trivial.
If $\mu_{N}\leq\bar{\mu}$, all experiments $(\hat{\mathbf{p}}(q),q)$ are robust to large deviations. This is in line with Proposition 4.3, which states that, fixing $\mu_{N}$, an experiment $(\hat{\mathbf{p}}(q),q)$ is robust to large deviations if and only if $q$ is sufficiently large. Moreover, when a trivial equilibrium exists, $V(\mu_{1},\mu_{1})=0$, so $(\hat{\mathbf{p}}(q),q)$ is a D1 pooling equilibrium if and only if $q\leq\bar{q}$. That is, the trivial D1 equilibrium coexists with a continuum of D1 pooling equilibria. Although the set of D1 pooling equilibrium outcomes is not closed, the trivial equilibrium outcome is a limit point in the sense that, as $q$ goes to zero, the players’ equilibrium payoffs and the joint distribution of the state and the receiver’s action converge to those in the trivial equilibrium.
5. Separating Equilibrium Under Single-Crossing
When the single-crossing property holds, there is no pooling in any D1 equilibrium. The following proposition shows that, when the cost of experiments is small, the D1 criterion selects a unique separating equilibrium outcome, and the receiver benefits from sender private information compared to the public-signal benchmark; when the cost of experiments is high, only the trivial equilibrium satisfies D1.
Proposition 5.1.
Let $C_{g}\leq\hat{K}(C_{b})$. There exists a unique D1 equilibrium outcome $\{\pi_{\theta}\}_{\theta\in\Theta}$ such that:
  1. (i)
    $\pi_{\theta}$ is the uninformative experiment if $V(\mu_{\theta},\mu_{\theta})=0$.
  2. (ii)
    $\pi_{\theta}$ is persuasive at belief $\mu_{\theta}$ if $V(\mu_{\theta},\mu_{\theta})>0$.
  3. (iii)
    The lowest sender type’s equilibrium payoff $v_{1}^{\star}=V(\mu_{1},\mu_{1})$. All sender types $\theta\geq 2$ such that $V(\mu_{\theta},\mu_{\theta})>0$ chooses the experiment $\pi_{\theta}=(p_{\theta},\mathbf{Q}(\mu_{\theta})p_{\theta})$, where $p_{\theta}$ is the unique solution of
    (5.1) $v^{\star}_{\theta}=$ $\max_{p}f((p,\mathbf{Q}(\mu_{\theta})p),\mu_{\theta})$
    $s.t.~{}f((p,\mathbf{Q}(\mu_{\theta})p),\mu_{\theta-1})\leq v_{\theta-1}^{\star}.$
Moreover, letting $U^{s}$ be the receiver’s payoff in a D1 equilibrium, $U^{s}\geq U^{ps}$.
The uniqueness result under single-crossing is reminiscent of Cho and Sobel,’s (1990) results despite differences in our settings. [^11] Cho and Sobel, (1990) assume that the signal space is the product of intervals, and higher sender types are more willing to send higher signals (with respect to the product order). In our game, since the cost of any fully revealing experiment is infinite, the signal space consists of $\Pi^{\circ}$ and the uninformative experiment, which is not compact. Moreover, the monotonicity assumption does not hold. The structure of the unique D1 equilibrium outcome in our game also differs from that in Cho and Sobel, (1990). Recall that in Cho and Sobel, (1990), the equilibrium signal is nondecreasing in the sender’s type, and some highest sender types may pool on the greatest signal. Recall from Lemma 3.2 that $V(\mu_{\theta},\mu_{\theta})$ is weakly increasing in the sender’s type. Therefore, conditions (i) and (ii) in Proposition 5.1 imply that a set of the lowest sender types may choose the uninformative experiment. When the cost of experiments is so high that $V(\mu_{N},\mu_{N})=0$, all sender types choose the uninformative experiment. Condition (iii) shows that the equilibrium outcome can be solved by induction, starting from the lowest sender type. The intuition is that no sender type can increase her payoff without violating a lower sender type’s incentive constraints. The single-crossing property allows us to solve a simplified maximization problem (5.1) which includes only the incentive constraint of the adjacent lower type.
Let us illustrate this in Figure 5.1. Suppose that there are two sender types $h>l$, and both of them have a positive payoff guarantee. That is, by choosing some experiment $\hat{\pi}_{\theta}$, the type $\theta$ sender gets payoff $V(\mu_{\theta},\mu_{l})>0$ even when the receiver believes that she is the low type. In a separating equilibrium, the receiver knows the sender’s type at the interim stage. Hence, the low-type sender gets no information rent. In all separating equilibria, she chooses the experiment $\hat{\pi}_{l}$ and receives her payoff guarantee, i.e., $v_{l}^{\star}=V(\mu_{l},\mu_{l})$.
[…]
Figure 5.1: The unique D1 equilibrium outcome under single-crossing
For the high-type sender, her choice of experiment $\pi_{h}=(p,q)$ must satisfy three constraints. First, it must give her at least her payoff guarantee, i.e., $f(\pi_{h},\mu_{h})\geq V(\mu_{h},\mu_{l})$. Second, it cannot be a profitable deviation for the low-type sender, i.e., $f(\pi_{h},\mu_{l})\leq v_{l}^{\star}$. Finally, it must be persuasive at the receiver’s interim belief $\mu_{h}$, i.e., $\frac{q}{p}\leq\mathbf{Q}(\mu_{h})$. The shaded area in Figure 5.1 shows the set of experiments $\Pi_{h}^{\star}$ satisfying these constraints. Among them, $\pi_{h}^{\star}$ uniquely maximizes the high-type sender’s payoff. The solid curve through $\pi_{h}^{\star}$ shows the high-type sender’s indifference curve.
Any experiment $\pi_{h}\in\Pi_{h}^{\star}$ is supported by a separating equilibrium, but the unique D1 equilibrium outcome is the one where $\pi_{h}=\pi_{h}^{\star}$. Consider a separating equilibrium where $\pi_{h}\neq\pi_{h}^{\star}$. We can find an experiment $\tilde{\pi}$, slightly below the experiment $\pi^{\star}_{h}$, which gives the high-type sender a higher payoff than $\pi_{h}$. That is, $\tilde{\pi}$ is a profitable deviation for the high type if the receiver’s interim belief is high, but deviating to $\tilde{\pi}$ is strictly equilibrium dominated for the low type. Hence, the D1 criterion requires that the receiver attributes this deviation to the high type. But given this off-path belief, $\tilde{\pi}$ is a profitable deviation for the high type. Therefore, the separating equilibrium does not satisfy D1. In contrast, when the high-type sender is prescribed to choose $\pi_{h}^{\star}$, any deviation that is profitable for the high-type sender is also profitable for the low-type sender. Hence, the D1 criterion is silent about the receiver’s interim belief.
The experiment $\pi_{h}^{\star}$ is Blackwell more informative than the experiment chosen by the high-type sender in the public-signal benchmark, which is marked by $\pi_{h}^{ps}$ in Figure 5.1. When there are more than two sender types, every sender type but the lowest chooses a weakly more Blackwell informative experiment than in the public-signal benchmark, in order to prevent lower types from mimicking. The receiver’s payoff is thus weakly higher in the D1 equilibrium than in the public-signal benchmark.
For the sender, the D1 equilibrium outcome is Pareto optimal among all separating equilibrium outcomes, that is, there does not exist a separating equilibrium outcome that gives all types of the sender weakly higher payoffs and at least one sender type a strictly higher payoff. This follows from Proposition 5.1, which states that it is not possible to increase any sender type’s equilibrium payoff without also raising the equilibrium payoff of the lower adjacent type. But in all separating equilibria, the lowest sender type has an equilibrium payoff equal to her payoff guarantee.
Corollary 5.2.
Let $C_{g}\leq\hat{K}(C_{b})$. The D1 equilibrium outcome is Pareto optimal among all separating equilibrium outcomes.
6. Discussions
6.1. Partial pooling equilibria
When the single-crossing property fails, partial pooling equilibria may also satisfy the D1 criterion. We present in this section some properties of these equilibria.
The nearby deviation argument still applies. Therefore, if an experiment $\pi=(p,q)$ is chosen by two sender types $h>l$, $p=\hat{\mathbf{p}}(q)$. Moreover, any sender type $m\in[l,h]$ must also choose $\pi$. Suppose that $\pi_{m}\neq\pi$. Since $f(\pi,\mu)-f(\pi_{m},\mu)$ is linear in $\mu$, it must be the case that all sender types are indifferent between $\pi$ and $\pi_{m}$. That is, $\pi_{m}$ is the experiment $\pi^{\prime}$ in Figure 4.3 where the indifference curves of all sender types through $\pi$ intersect again. $\pi_{m}$ is persuasive at belief $\mu_{m}$, since it is chosen only by the type $m$ sender. Therefore, there exists a deviation $\tilde{\pi}$ that is persuasive at belief $\mu_{h}$, gives the type $h$ sender a higher payoff than $\pi$ but yields a strictly lower payoff than $\pi$ (hence is dominated by equilibrium play) for all sender types below $h$. By the D1 criterion, the receiver should assign at least probability $\mu_{h}$ to the high state after seeing this deviation, but this makes $\tilde{\pi}$ a profitable deviation for the type $h$ sender. Finally, as the sender’s payoff from choosing $(\hat{\mathbf{p}}(q),q)$ increases in $q$ for all types, in any partial pooling D1 equilibrium, there is a single interval of sender types $[\underline{\theta},\bar{\theta}]$ who pool on the same experiment, while all other sender types choose distinct experiments.
If $V(\mu_{1},\mu_{1})=0$, no experiment that is persuasive at belief $\mu_{1}$ gives the lowest sender type a positive payoff, but any experiment $(\hat{\mathbf{p}}(q),q)$ gives her a positive payoff. Therefore, the lowest sender type is one of the pooling types. That is, any D1 partial pooling equilibrium features a set of the lowest sender types who choose the same experiment when the cost of experiments is high.
The large deviation argument also applies to partial pooling equilibria. That is, for all experiments $\tilde{\pi}$ that are persuasive at belief $\mu_{\bar{\theta}}$, $f(\tilde{\pi},\mu_{\bar{\theta}})>f(\pi,\mu_{\bar{\theta}})$ implies $f(\tilde{\pi},\mu_{\theta})>f(\pi,\mu_{\theta})$ for all $\theta<\bar{\theta}$. By Proposition 4.3, this is equivalent to a lower bound $\hat{q}_{\bar{\theta}}$ on $q$, but this lower bound depends on the highest pooling type $\bar{\theta}$.
6.2. Limiting cases
In the classic signaling model, there is a discontinuity as the sender’s private information becomes degenerate or the differential cost of sending signals diminishes. These discontinuities persist in our game, where the signal is a public experiment that reveals information about the payoff relevant state of the world.
When there is no sender private information, the no-signal benchmark with common prior $\mu_{h}$ has a unique equilibrium outcome $\pi^{ns}$. Now suppose that the sender has slight private information such that her prior is $\mu_{l}<\mu_{h}$ with an infinitesimal probability. One of the following happens depending on the cost of experiments. If $C_{g}\leq\hat{K}(C_{b})$, the game has a unique separating equilibrium, where the high-type sender chooses a more informative experiment than $\pi^{ns}$. The separating outcome is independent of the probability of the low type sender, hence it does not converge to $\pi^{ns}$ as the probability goes to zero. This is akin to the discontinuity observed in the classic signaling model. If $C_{g}>\hat{K}(C_{b})$, the game has a continuum of pooling equilibria, and all pooling equilibrium outcomes satisfy $p=\hat{\mathbf{p}}(q)$ and are therefore separated from $\pi^{ns}$. Therefore, although the failure of single-crossing leads to multiplicity of equilibrium outcomes, none of them converges to the equilibrium outcome of the no-signal benchmark as the sender’s private information becomes degenerate.
When experiments are costless, the sender’s payoff from choosing a persuasive experiment simply equals the probability of the good outcome. All equilibria are persuasive and pooling, and there exists a continuum of D1 pooling equilibria where the sender chooses the experiment $(1,q)$ for some $q\in[\mathbf{Q}(\mu_{1}),\mathbf{Q}(\mu_{0})]$. That is, the bad outcome reveals the bad state, and the receiver’s posterior belief after seeing the good outcome is weakly higher than the decision threshold $\bar{\beta}$. Now consider a sequence of costly persuasion games with diminishing costs such that $\frac{C_{g}^{n}}{C_{b}^{n}}$ is uniformly bounded, where $C_{g}^{n}$ and $C_{b}^{n}$ denote the costs of good news and bad news in the $n$’th game, respectively. Recall from Proposition 3.1 that $\lim_{C_{b}\downarrow 0}\hat{K}^{\prime}(C_{b})=\infty$. Therefore, as the costs diminish, the game has a unique separating equilibrium outcome. In the limit, it converges to a strategy profile where the type $\theta$ sender chooses the experiment $(1,\mathbf{Q}(\mu_{\theta}))$.
6.3. Microfoundation of the log-likelihood ratio cost
The log-likelihood ratio cost of experiments can be microfounded by a Wald sequential sampling problem. The sender acquires noisy, binary signals about the state of the world before making an irreversible decision to stop. An experiment in the persuasion game is equivalent to a threshold stopping rule in the sampling problem, that is, the sender stops acquiring signals at the first instance her belief reaches some thresholds. Equating the cost of an experiment with the expected cost of acquiring signals in the sampling problem yields the log-likelihood ratio costs.
Consider the following stopping problem in discrete time. [^12] A continuous time version of the model is a special case of that in Morris and Strack, (2019). The sender observes a Brownian motion with a state-dependent drift term, and her flow cost $c(p_{t})$ is a function of her instantaneous belief $p_{t}$ on the good state. Motivated by our application, it is straightforward to set $c(p_{t})=p_{t}c_{g}+(1-p_{t})c_{b}$. That is, conditional on the state $\omega$, the flow cost of experimenting is a constant $c_{\omega}$. Then the log-likelihood ratio cost function is derived using Theorem 1 of Morris and Strack, (2019). There is a binary state of the world $\omega\in\{G,B\}$, and a sequence of binary signals $(s_{n})_{n=1}^{\infty}$ such that each $s_{n}\in\{g,b\}$. Let $\mu_{0}$ be the probability of the good state, and conditional on the state, signals are distributed iid such that $\mathbb{P}(s_{n}=g|\omega=G)=\mathbb{P}(s_{n}=b|\omega=B)=\alpha>\frac{1}{2}$. Both the state of the world and the signals are realized at the outset of the game and are not observed by the sender.
We model the sender’s information acquisition as follows. At any history $h_{n}=(s_{1},s_{2},\dots,s_{n})$ of the game, that is, the sender has acquired signals $s_{1},s_{2},\dots,s_{n}$ and has not yet stopped, the sender chooses between acquiring an additional signal $s_{n+1}$ and irreversibly stopping signal acquisition. The cost of acquiring the signal $s_{n+1}$ is a random variable that equals $c_{g}$ if $s_{n+1}=g$ and $c_{b}$ if $s_{n+1}=b$. Hence, the total cost of signals if the sender stops at history $h_{n}$ is $c_{g}n_{g}(h_{n})+c_{b}n_{b}(h_{n})$, where $n_{g}(h_{n})$ and $n_{b}(h_{n})$ denote the number of good and bad signals in $h_{n}$, respectively.
A strategy of the sender is a stopping time adapted to the natural filtration generated by histories. Specifically, consider the following threshold strategy $\tau$ of the sender: she stops at the first history where the difference between the number of good and bad signals equals some threshold values $\underline{n}<0$ or $\bar{n}>0$. Notice that the sender stops in finite time with probability one. Hence, interpreting the event $n_{g}(h_{\tau})-n_{b}(h_{\tau})=\bar{n}$ as the good outcome and the event $n_{g}(h_{\tau})-n_{b}(h_{\tau})=\underline{n}$ as the bad outcome, the strategy $\tau$ is equivalent to an experiment $\pi=(p,q)$ in the persuasion game, where $p$ and $q$ are the probabilities of the good outcome conditional on the good and bad state, respectively.
The sender’s posterior belief on the good state induced by the stopping rule $\tau$ is a random variable $\hat{\mu}$, which equals $\mu_{\bar{n}}:=\left[1+\frac{1-\mu_{0}}{\mu_{0}}\left(\frac{1-\alpha}{\alpha}\right)^{\bar{n}}\right]^{-1}$ following the good outcome, or $\mu_{\underline{n}}:=\left[1+\frac{1-\mu_{0}}{\mu_{0}}\left(\frac{1-\alpha}{\alpha}\right)^{\underline{n}}\right]^{-1}$ following the bad outcome. This is the same as the posterior belief induced by the experiment $\pi$.
Proposition 6.1.
The expected cost of implementing the strategy $\tau$, $\mathbb{E}[c_{g}n_{g}(h_{\tau})+c_{b}n_{b}(h_{\tau})],$ equals $\mathbb{E}[H(\mu_{0})-H(\hat{\mu})]$, where
$H(\mu)=\frac{1}{2\alpha-1}\left[\ln\left(\frac{\alpha}{1-\alpha}\right)\right]^{-1}\left[\bar{c}_{g}\mu\ln\left(\frac{1-\mu}{\mu}\right)+\bar{c_{b}}(1-\mu)\ln\left(\frac{\mu}{1-\mu}\right)\right],$
and $\bar{c}_{g}=\alpha c_{g}+(1-\alpha)c_{b}$, $\bar{c}_{b}=(1-\alpha)c_{g}+\alpha c_{b}$.
Hence, letting $C_{g}=\frac{1}{2\alpha-1}\left[\ln\left(\frac{\alpha}{1-\alpha}\right)\right]^{-1}\bar{c}_{g}$ and $C_{b}=\frac{1}{2\alpha-1}\left[\ln\left(\frac{\alpha}{1-\alpha}\right)\right]^{-1}\bar{c_{b}}$, the expected cost of implementing the strategy $\tau$ equals the log-likelihood ratio cost of the experiment $\pi$.
6.4. Shannon entropy cost of experiments
Our result that the single-crossing property may fail in a costly persuasion game with a partially informed sender is not specific to the log-likelihood ratio cost of experiments we use. The following proposition shows a similar result in a game with the Shannon entropy cost of experiments. That is, $c(\pi|\mu)=\mathbb{E}[H(\mu)-H(\hat{\mu})]$, where $H(\mu)=-C\left[\mu\ln\mu+(1-\mu)\ln(1-\mu)\right]$, and $C>0$.
Proposition 6.2.
The single-crossing property fails in the persuasion game with the Shannon entropy cost of experiments.
  1. (i)
    There exists $\tilde{\mathbf{p}}:(0,1)\to(0,1)$ such that $\tilde{\mathbf{p}}(q)>q$, and $MRS(\pi|\mu)$ is decreasing in $\mu$ if $p\leq\tilde{\mathbf{p}}(q)$, and $MRS(\pi|\mu)$ first increases and then decreases in $\mu$ if $p>\tilde{\mathbf{p}}(q)$;
  2. (ii)
    For any sender types $i<j$, there exists $\tilde{\mathbf{p}}_{i,j}:(0,1)\to(0,1)$ such that $\tilde{\mathbf{p}}_{i,j}(q)>\tilde{\mathbf{p}}(q)$, and $MRS(\pi|\mu_{i})>MRS(\pi|\mu_{j})$ if $p<\tilde{\mathbf{p}}_{i,j}(q)$, $MRS(\pi|\mu_{i})=MRS(\pi|\mu_{j})$ if $p=\tilde{\mathbf{p}}_{i,j}(q)$, and $MRS(\pi|\mu_{i})<MRS(\pi|\mu_{j})$ if $p>\tilde{\mathbf{p}}_{i,j}(q)$.
Part (i) of Proposition 6.2 shows that the single-crossing property fails at experiments that are very informative (i.e., $p>\tilde{\mathbf{p}}(q)$). In fact, the double-crossing property (Chen et al.,, 2022) is satisfied. As part (ii) shows, given any two sender types, there exists a locus of experiments where the indifference curves of the two sender types are tangent, and the high-type sender’s indifference curve is more convex. Therefore, the D1 criterion may select an equilibrium that features pairwise pooling where each experiment is chosen by up to two sender types.
7. Conclusions
We study a persuasion game with costly experiments and a partially informed sender. At the outset of the game, the sender privately observes a noisy signal about the state of the world. She then chooses a costly experiment which generates public information to the receiver. The receiver can infer the state of the world from the choice of experiment as well as the outcome of the experiment.
The equilibrium outcome crucially depends on the cost of experiments. We focus on the log-likelihood ratio cost where the costs of drawing good news and bad news can differ. If bad news is more costly than good news, there exists a unique separating equilibrium outcome. The receiver fully learns the sender’s private signal from her choice of experiment, and the sender chooses a more informative experiment than what she would have chosen in the public-signal benchmark. Therefore, the receiver is unambiguously better off with sender private information. On the other hand, if good news is more costly than bad news, this is a signaling game where the single-crossing property fails. There may exist pooling equilibria, that is, the receiver learns nothing from the sender’s choice of experiment, and in some pooling equilibria, the receiver’s payoff is strictly lower than that in the benchmark with no sender private information.
These results have practical implications for different real-life applications. In drug approval, information asymmetry at the preclinical stage incentivizes pharmaceutical companies to run more informative clinical trials. Hence, it improves public welfare. In startup funding, however, an entrepreneur’s private knowledge about a proprietary technology can hurt investors. Therefore, it is beneficial to solicit opinions from external experts.
Appendix A Auxiliary Results and Proofs
A.1. Equilibrium outcome of the symmetric information benchmark
Lemma A.1.
In the symmetric information benchmark, $V(\mu_{s},\mu_{r})>0$ if and only if $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})>0$, where
$\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})=\mu_{s}+(1-\mu_{s})\mathbf{Q}(\mu_{r})+$ $\mu_{s}C_{g}\left[\ln\mathbf{Q}(\mu_{r})+1-\mathbf{Q}(\mu_{r})\right]$
$\quad-(1-\mu_{s})C_{b}\left[\mathbf{Q}(\mu_{r})\ln\mathbf{Q}(\mu_{r})+1-\mathbf{Q}(\mu_{r})\right].$
If $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})>0$, there exists a unique equilibrium outcome which is persuasive at belief $\mu_{r}$. If $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})\leq 0$, the unique equilibrium outcome is the uninformative experiment.
Proof.
Notice that $f((p,\mathbf{Q}(\mu_{r})p),\mu_{s})$ is continuous in $p$ on $[0,1)$ and concave in $p$. Therefore, the sender’s maximization problem (3.1) is solved by the first order condition. To simplify notation, we will denote briefly the derivative $\frac{d}{dp}f((p,\mathbf{Q}(\mu_{r})p),\mu_{s})$ by $f^{\prime}(p)$, and $\mathbf{Q}(\mu_{r})$ by $Q$. Notice that
(A.1) $f^{\prime}(p)=\mu_{s}M(p,Qp)+(1-\mu_{s})N(p,Qp),$
where
$M(p,q):=1+C_{g}\left[\frac{p-q}{p(1-q)}-\ln\frac{p(1-q)}{q(1-p)}\right],$
$N(p,q):=\frac{q}{p}\left(1-C_{b}\left[\frac{p-q}{(1-p)q}-\ln\frac{p(1-q)}{q(1-p)}\right]\right).$
As $p\uparrow 1$, $f^{\prime}(p)$ goes to $\infty$- ∞. Therefore, the first order condition $f^{\prime}(p)=0$ has a unique interior solution $\hat{p}\in(0,1)$ if and only if
$\lim_{p\downarrow 0}f^{\prime}(p)=\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})>0.$
If $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})>0$, there exists a unique equilibrium where the sender chooses the experiment $\hat{\pi}=(\hat{p},Q\hat{p})$, and $V(\mu_{s},\mu_{r})>0$. If $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})\leq 0$, $f^{\prime}(p)<0$ for all $p\in(0,1)$. Hence, the sender’s maximization problem has the corner solution $\hat{p}=0$, and $V(\mu_{s},\mu_{r})=0$. In this case, the sender can obtain zero payoff if and only if she chooses the uninformative experiment, thus the uninformative experiment is the unique equilibrium outcome. ∎
Remarks.
Notice that for all $Q\in(0,1)$, $\ln Q+1-Q<0$, and $Q\ln Q+1-Q>0$. Therefore, $\mathbf{F}$ is a linear and decreasing function of $C_{g}$ and $C_{b}$. Letting
(A.2) $\overline{C_{g}}(\mu_{r})=\frac{\mathbf{Q}(\mu_{r})}{\mathbf{Q}(\mu_{r})\ln\mathbf{Q}(\mu_{r})+1-\mathbf{Q}(\mu_{r})},\qquad\overline{C_{b}}(\mu_{r})=-\frac{1}{\ln\mathbf{Q}(\mu_{r})+1-\mathbf{Q}(\mu_{r})},$
we notice that $\mathbf{F}(\overline{C_{g}}(\mu_{r}),\overline{C_{b}}(\mu_{r}),\mu_{s},\mu_{r})=0$ for all $\mu_{s},\mu_{r}\in(0,1)$. Therefore, persuasion is possible if and only if a weighted average of $C_{g}$ and $C_{b}$ is below a threshold. This corresponds to the region below a straight line through $(\overline{C_{b}}(\mu_{r}),\overline{C_{g}}(\mu_{r}))$ in the space of the cost parameters $(C_{b},C_{g})$, as is shown in Figure 3.1. As the sender’s prior $\mu_{s}$ increase, the slope of this line
$\frac{1-\mu_{s}}{\mu_{s}}\frac{\mathbf{Q}(\mu_{r})\ln\mathbf{Q}(\mu_{r})+1-\mathbf{Q}(\mu_{r})}{\ln\mathbf{Q}(\mu_{r})+1-\mathbf{Q}(\mu_{r})}$
decreases in absolute value. As $\mu_{s}\to 0$, the slope goes to negative infinity, and as $\mu_{s}\to 1$, the slope converges to zero.
A.2. Proof of Proposition 3.1
Proof.
Fix the receiver’s prior $\mu_{r}$. Consider first the case $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})>0$. That is, there is a unique persuasion equilibrium outcome $\hat{\pi}=(\hat{p},Q\hat{p})$, where $\hat{p}\in(0,1)$ is solved by the first order condition $f^{\prime}(p)=0$. By (A.1), $M(\hat{p},Q\hat{p})\geq 0\geq N(\hat{p},Q\hat{p})$ or $M(\hat{p},Q\hat{p})\leq 0\leq N(\hat{p},Q\hat{p})$. Hence, $\hat{p}$ is independent of $\mu_{s}$, that is,
$\frac{\partial}{\partial\mu_{s}}f^{\prime}(\hat{p})=M(\hat{p},Q\hat{p})-N(\hat{p},Q\hat{p})=0$
if and only if $M(\hat{p},Q\hat{p})=N(\hat{p},Q\hat{p})=0$.
Fixing $C_{b}>0$, $M(\hat{p},Q\hat{p})=N(\hat{p},Q\hat{p})=0$ defines a system of equations of $\hat{p}$ and $C_{g}$. It has a unique solution
(A.3) $\hat{p}=1-\frac{1-Q}{Q}\frac{1}{\hat{x}(C_{b})},$
(A.4) $\quad C_{g}=\hat{K}(C_{b}):=\left(\frac{\hat{x}(C_{b})^{2}}{1+\hat{x}(C_{b})}-\frac{1}{C_{b}}\right)^{-1},$
where $x=\hat{x}(C_{b})>0$ solves
(A.5) $x-\ln(1+x)=\frac{1}{C_{b}}.$
Notice that the left-hand side of (A.5) is increasing in $x$ on $\mathbb{R}_{+}$, and it equals zero at $x=0$ and goes to infinity as $x\to\infty$. Hence, $\hat{x}(C_{b})$ is well defined. If $C_{g}>\hat{K}(C_{b})$, $M(\hat{p},Q\hat{p})<0<N(\hat{p},Q\hat{p})$, hence $\frac{\partial}{\partial\mu_{s}}f^{\prime}(\hat{p})<0$, and $\hat{p}$ is strictly decreasing in $\mu_{s}$. If $C_{g}<\hat{K}(C_{b})$, $M(\hat{p},Q\hat{p})>0>N(\hat{p},Q\hat{p})$, hence $\frac{\partial}{\partial\mu_{s}}f^{\prime}(\hat{p})>0$, and $\hat{p}$ is strictly increasing in $\mu_{s}$.
Substituting $C_{b}=\overline{C_{b}}(\mu_{r})$ using (A.2) in (A.5), we solve that
(A.6) $\hat{x}\left(\overline{C_{b}}(\mu_{r})\right)=\frac{1-\mathbf{Q}(\mu_{r})}{\mathbf{Q}(\mu_{r})}.$
Substituting (A.6) in (A.4), we can verify that $\hat{K}\left(\overline{C_{b}}(\mu_{r})\right)=\overline{C_{g}}(\mu_{r})$ for all $\mu_{r}\in(0,1)$. Recall that when $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})\leq 0$, the sender’s optimization problem has the corner solution $\hat{p}=0$. Hence, combining the two cases, we conclude that $\hat{p}$ is independent of $\mu_{s}$ if $C_{g}=\hat{K}(C_{b})$, weakly increasing in $\mu_{s}$ if $C_{g}<\hat{K}(C_{g})$, and weakly decreasing in $\mu_{s}$ if $C_{g}>\hat{K}(C_{b})$. Equivalently, the equilibrium outcome $\hat{\pi}$ is independent of $\mu_{s}$ if $C_{g}=\hat{K}(C_{b})$, more Blackwell informative as $\mu_{s}$ increases if $C_{g}<\hat{K}(C_{g})$, and less Blackwell informative as $\mu_{s}$ increases if $C_{g}>\hat{K}(C_{b})$.
Notice that $\hat{x}$ is a smooth function of $C_{b}$. Therefore, $\hat{K}$ is a smooth function of $C_{b}$. We are left to check that $\hat{K}$ is increasing and concave, and that $\lim_{C_{b}\downarrow 0}\hat{K}^{\prime}(C_{b})=\infty$. By (A.5),
$\hat{x}^{\prime}(C_{b})=-\frac{1}{C_{b}^{2}}\left(1-\frac{1}{1+\hat{x}(C_{b})}\right)^{-1}.$
Hence, taking derivative of (A.4),
(A.7) $\hat{K}^{\prime}(C_{b})=\left(\frac{\hat{K}(C_{b})}{C_{b}}\right)^{2}\frac{1}{1+\hat{x}(C_{b})}>0.$
As $C_{b}\downarrow 0$, applying l’Hôpital’s rule on (A.7) yields
$\lim_{C_{b}=0}\frac{1}{\hat{K}^{\prime}(C_{b})}=\lim_{C_{b}=0}\frac{1}{1+\hat{x}(C_{b})}=0.$
Therefore, $\lim_{C_{b}\downarrow 0}\hat{K}^{\prime}(C_{b})=\infty$.
Notice that $[(1+\hat{x}(C_{b}))\hat{K}^{\prime}(C_{b})]^{\prime}=\hat{x}^{\prime}(C_{b})\hat{K}^{\prime}(C_{b})+(1+\hat{x}(C_{b}))\hat{K}^{\prime\prime}(C_{b})$. Hence,
$\hat{K}^{\prime\prime}(C_{b})$ $=\frac{[(1+\hat{x}(C_{b}))\hat{K}^{\prime}(C_{b})]^{\prime}-\hat{x}^{\prime}(C_{b})\hat{K}^{\prime}(C_{b})}{1+\hat{x}(C_{b})}$
(A.8) $=\frac{\hat{K}(C_{b})^{2}}{(1+\hat{x}(C_{b}))C_{b}^{3}}\left[2\left(\frac{\hat{K}(C_{b})}{C_{b}(1+\hat{x}(C_{b})){}}-1\right)+\frac{1}{C_{b}\hat{x}(C_{b})}\right].$
Substituting $\hat{K}(C_{b})$ using (A.4) and $\frac{1}{C_{b}}$ using (A.5), and to shorten notation, letting $x=\hat{x}(C_{b})$, the bracket in (A.8) becomes
(A.9) $\frac{3x^{2}-(x^{2}+2x)\ln(1+x)-(1+x)(\ln(1+x))^{2}}{(1+x)\ln(1+x)-x}.$
The denominator of (A.9) is positive for all $x>0$. Taking third order derivative of the numerator of (A.9) with respect to $x$ yields
$\frac{-2x(2+x)+2(1+x)\ln(1+x)}{(1+x)^{3}}<\frac{2[\ln(1+x)-x]}{(1+x)^{2}},$
which is negative for all $x>0$. Notice that evaluated at $x=0$, the numerator of (A.9) and its derivatives up to the third order are zero. Therefore, the numerator of (A.9) is negative for all $x>0$. Hence, (A.9) is negative, and $\hat{K}^{\prime\prime}(C_{b})<0$. That is, $\hat{K}$ is increasing and concave. ∎
A.3. Proof of Lemma 3.2
Proof.
Let $\bar{U}(\mu)$ be the receiver’s equilibrium payoff in the symmetric information benchmark with common prior $\mu_{s}=\mu_{r}=\mu$. We are to show that $\bar{U}$ is piecewise linear and weakly convex, hence
$U^{ns}=\bar{U}(\mu_{0})\leq\mathbb{E}[\bar{U}(\mu_{\theta})]=U^{ps}.$
Given $\mu<\bar{\beta}$, in the symmetric information game with common prior $\mu$, the sender solves
(A.10) $V(\mu,\mu)=\max_{x\leq\mu}\frac{\mu-x}{\bar{\beta}-x}-\left(H(\mu)-\frac{\mu-x}{\bar{\beta}-x}H(\bar{\beta})-\frac{\bar{\beta}-\mu}{\bar{\beta}-x}H(x)\right).$
The first term in (A.10) is the probability of the good outcome if the sender chooses an experiment that splits the receiver’s belief to $\bar{\beta}$ and $x$, and the bracket in (A.10) is the associated cost. The sender’s problem (A.10) admits a unique solution $x^{\star}$. If $x^{\star}<\mu$, in the equilibrium, the sender chooses an experiment that splits the receiver’s belief to $x^{\star}$ and $\bar{\beta}$. If $x^{\star}=\mu$, persuasion is not possible–the sender chooses an uninformative experiment in any equilibrium.
The first order condition of (A.10) is
(A.11) $\frac{\bar{\beta}-\mu}{(\bar{\beta}-x)^{2}}\left[1+H(\bar{\beta})-H(x)\right]-\frac{\bar{\beta}-\mu}{\bar{\beta}-x}H^{\prime}(x)=0.$
That is,
[…]
Taking derivative of the left-hand side of (A.12) with respect to $x$ yields $H^{\prime\prime}(x)(\bar{\beta}-x)<0$. Moreover, as $x\to 0$, the left-hand side of (A.12) goes to $\infty$. Hence, (A.12) has a unique solution $\underline{\beta}<\bar{\beta}$ which is independent of $\mu$. If $\underline{\beta}<\mu<\bar{\beta}$, $x^{\star}=\underline{\beta}$ solves the sender’s problem; if $\mu<\underline{\beta}$, the sender’s problem has the corner solution $x^{\star}=\mu$.
To summarize, if $\underline{\beta}<\mu<\bar{\beta}$, the unique equilibrium outcome splits the players’ belief to $\underline{\beta}$ and $\bar{\beta}$, and the receiver’s expected payoff $\bar{U}(\mu)=\frac{\mu-\underline{\beta}}{\bar{\beta}-\underline{\beta}}\bar{U}(\bar{\beta})$; if $\mu\leq\underline{\beta}$, persuasion is not possible, and the receiver’s payoff $\bar{U}(\mu)=0$. Therefore, we conclude that $\bar{U}$ is piecewise linear and weakly convex.
Applying envelope theorem to (A.10) and using the fact that (A.11) holds at $x=\underline{\beta}$,
$\frac{d}{d\mu}V(\mu,\mu)=\frac{1}{\bar{\beta}-\underline{\beta}}\left[1+H(\bar{\beta})-H(\underline{\beta})\right]-H^{\prime}(\mu)=H^{\prime}(\underline{\beta})-H^{\prime}(\mu)>0$
if $\underline{\beta}<\mu<\bar{\beta}$. If $\mu\leq\underline{\beta}$, $V(\mu,\mu)=0$. Hence, the sender’s equilibrium payoff $V(\mu,\mu)$ is weakly increasing in $\mu$. ∎
A.4. Proof of Proposition 4.1
Proof.
Let $\mu\in(0,1)$ and $\pi=(p,q)\in\Pi^{\circ}$. Observe that
(A.13) $\frac{\partial f(\pi,\mu)}{\partial p}=-\underbrace{C_{b}\frac{p-q}{p(1-p)}}_{=:A_{1}(p,q)}-\mu\underbrace{\left[-1+C_{g}\ln\frac{p(1-q)}{(1-p)q}-C_{b}\frac{p-q}{p(1-p)}\right]}_{=:A_{2}(p,q)},$
(A.14) $\frac{\partial f(\pi,\mu)}{\partial q}=\underbrace{1+C_{b}\ln\frac{p(1-q)}{(1-p)q}}_{=:A_{3}(p,q)}+\mu\underbrace{\left[-1+C_{g}\frac{p-q}{q(1-q)}-C_{b}\ln\frac{p(1-q)}{(1-p)q}\right]}_{=:A_{4}(p,q)}.$
Hence,
$MRS(\pi|\mu)=-\frac{\partial f(\pi,\mu)/\partial p}{\partial f(\pi,\mu)/\partial q}=\frac{A_{1}(p,q)+A_{2}(p,q)\mu}{A_{3}(p,q)+A_{4}(p,q)\mu},$
and
$\frac{\partial}{\partial\mu}MRS(\pi|\mu)=\frac{\Delta(p,q)}{\big{[}A_{3}(p,q)+A_{4}(p,q)\mu\big{]}^{2}},$
where
$\Delta(p,q)$ $=A_{2}(p,q)A_{3}(p,q)-A_{1}(p,q)A_{4}(p,q)$
$=(C_{g}-C_{b})\ln\frac{p(1-q)}{(1-p)q}+C_{g}C_{b}\left[\left(\ln\frac{p(1-q)}{(1-p)q}\right)^{2}-\frac{(p-q)^{2}}{pq(1-p)(1-q)}\right]-1.$
Notice that $\Delta(p,q)$ is independent of $\mu$. Therefore, the marginal rate of substitution at any given experiment is monotonic in the sender’s prior belief, and whether it is increasing or decreasing is determined by the sign of $\Delta(p,q)$.
Define a change of variable
(A.15) $t=\frac{(1-p)q}{p(1-q)}.$
$t\in(0,1)$ for all $(p,q)\in\Pi^{\circ}$. We can rewrite $\Delta$ as a function of just $t$. To simplify notation, we use the same letter $\Delta$ to denote this function
(A.16) $\Delta(t)=-(C_{g}-C_{b})\ln t+C_{g}C_{b}\left[(\ln t)^{2}-\frac{(1-t)^{2}}{t}\right]-1.$
Taking derivative of (A.16),
(A.17) $\Delta^{\prime}(t)=-\frac{C_{g}-C_{b}}{t}+\frac{C_{g}C_{b}}{t}\left(2\ln t-t+\frac{1}{t}\right).$
Hence, $\Delta^{\prime}(t)=0$ if and only if
(A.18) $2\ln t-t+\frac{1}{t}=\frac{1}{C_{b}}-\frac{1}{C_{g}}.$
The left-hand side of (A.18) is positive for all $t\in(0,1)$ and decreasing in $t$. As $t\to 0$, it goes to infinity, and at $t=1$, it equals zero. Therefore, if $C_{g}\leq C_{b}$, $\Delta^{\prime}(t)>0$ for all $t\in(0,1)$, so $\Delta(t)<\Delta(1)=-1$ for all $t\in(0,1)$. That is, the marginal rate of substitution at any experiment is decreasing in the sender’s prior belief, so the single-crossing property holds.
If $C_{g}>C_{b}$, (A.18) has a unique solution $t^{\star}\in(0,1)$. $\Delta(t)$ is single-peaked and obtains its maximum at $t=t^{\star}$. Therefore, the single-crossing property holds if and only if $\Delta(t^{\star})\leq 0$. Notice that $\Delta(1)=-1$ and $\Delta^{\prime}(1)<0$, so $\Delta(t^{\star})>-1$. By envelope theorem,
(A.19) $\frac{\partial\Delta(t^{\star})}{\partial C_{g}}=-\ln t^{\star}+C_{b}\left[(\ln t^{\star})^{2}-\frac{(t^{\star}-1)^{2}}{t^{\star}}\right]\geq\frac{\Delta(t^{\star})+1}{C_{g}}>0.$
Notice that $y^{\prime}(x)=\frac{y(x)+1}{x}$ solves a linear function $y$ of $x$. Therefore, (A.19) implies that $\Delta(t^{\star})$ increases in $C_{g}$ at least as fast as a linear function. Moreover, given any $C_{b}>0$, as $C_{g}\downarrow C_{b}$, $t^{\star}\to 1$, and $\Delta(t^{\star})<0$. Therefore, $\Delta(t^{\star})=0$ has a unique solution $C_{g}=\tilde{K}(C_{b})$.
We now show that $\tilde{K}$ is indeed the same function $\hat{K}$ in Proposition 3.1. Fix $\mu_{r}\in(0,\bar{\beta})$ and let $Q=\mathbf{Q}(\mu_{r})$. Substituting $C_{g}=\overline{C_{g}}(\mu_{r})$ and $C_{b}=\overline{C_{b}}(\mu_{r})$ using (A.2) in (A.17),
(A.20) $\Delta^{\prime}(t)=\frac{1}{t}\frac{2Q\ln Q+1-Q^{2}-Q\left(2\ln t-t+\frac{1}{t}\right)}{(\ln Q+1-Q)(Q\ln Q+1-Q)}.$
It is easy to verify that (A.20) has a unique zero $t^{\star}=Q$. Evaluating (A.16) at $t^{\star}=Q$, we have
$\Delta(t^{\star})=\frac{Q(\ln Q)^{2}+(1-Q^{2})\ln Q+(1-Q)^{2}}{(\ln Q+1-Q)(Q\ln Q+1-Q)}-1=0$
That is, $\overline{C_{g}}(\mu_{r})=\tilde{K}\left(\overline{C_{b}}(\mu_{r})\right)$ for all $\mu_{r}\in(0,\bar{\beta})$. Recall that $\overline{C_{g}}(\mu_{r})=\hat{K}\left(\overline{C_{b}}(\mu_{r})\right)$ for all $\mu_{r}\in(0,\bar{\beta})$, and $\overline{C_{b}}:(0,\bar{\beta})\to(0,\infty)$ is injective. Hence, $\tilde{K}(C_{b})=\hat{K}(C_{b})$ for all $C_{b}>0$.
If $C_{g}\leq\hat{K}(C_{b})$, $\Delta(t)\leq\Delta(t^{\star})\leq 0$ for all $t\in(0,1)$. That is, $\frac{\partial}{\partial\mu}MRS(\pi|\mu)\leq 0$ for all $\pi\in\Pi^{\circ}$, and the single-crossing property holds. If $C_{g}>\hat{K}(C_{b})$, $\Delta(t^{\star})>0$. Therefore, $\Delta(t)$ has two zeros $\hat{t}$ and $\check{t}$ such that $01>\hat{t}>t^{\star}>\check{t}>$ 0, and
$\Delta(t)\left\{\begin{array}[]{ll}<0&\text{if $t\in(0,\check{t})\cup(\hat{t},1)$}\\ =0&\text{if $t\in\{\check{t},\hat{t}\}$}\\ >0&\text{if $t\in(\check{t},\hat{t})$}\end{array}\right..$
This is equivalent to the desired result (4.2), where
(A.21) $\hat{\mathbf{p}}(q)=\frac{q}{q+\hat{t}(1-q)},\qquad\check{\mathbf{p}}(q)=\frac{q}{q+\check{t}(1-q)}.$
Notice that
$\frac{\partial\Delta(t)}{\partial C_{b}}=\ln t+C_{g}\left[(\ln t)^{2}-\frac{(1-t)^{2}}{t}\right]<0$
for all $t\in(0,1)$. Therefore, $\hat{t}$ is decreasing in $C_{b}$, and $\check{t}$ is increasing in $C_{b}$. Equivalently, $\hat{\mathbf{p}}(q)$ is increasing in $C_{b}$, and $\check{\mathbf{p}}(q)$ is decreasing in $C_{b}$ for all $q\in(0,1)$. On the other hand,
$\frac{\partial\Delta(t)}{\partial C_{g}}=-\ln t+C_{b}\left[(\ln t)^{2}-\frac{(1-t)^{2}}{t}\right]=\frac{\Delta(t)+1-C_{b}\ln t}{C_{g}}.$
Hence, at $\hat{t}$ and $\check{t}$,
$\frac{\partial\Delta(t)}{\partial C_{g}}=\frac{1-C_{b}\ln t}{C_{g}}>0.$
Therefore, $\hat{t}$ is increasing in $C_{g}$, and $\check{t}$ is decreasing in $C_{b}$. Equivalently, $\hat{\mathbf{p}}(q)$ is decreasing in $C_{g}$, and $\check{\mathbf{p}}(q)$ is increasing in $C_{g}$. ∎
A.5. Results relating to Proposition 4.1
Lemma A.2.
Let $C_{g}>\hat{K}(C_{b})$. $M(p,q),N(p,q)>0$ if $q<p\leq\hat{\mathbf{p}}(q)$, and $M(p,q),N(p,q)<0$ if $\check{\mathbf{p}}(q)\leq p<1$.
Proof.
Using the change of variable (A.15), we can rewrite $M(p,q)$ and $N(p,q)$ as
$M(p,q)=1+C_{g}(1-t+\ln t)=:m(t)$
$N(p,q)=\frac{q}{p}\left[1-C_{b}\left(\frac{1}{t}-1+\ln t\right)\right]=:\frac{q}{p}\cdot n(t).$
Both $m(t)$ and $n(t)$ are increasing in $t$. Hence, it is sufficient to show that $m(\hat{t}),n(\hat{t})>0$ and $m(\check{t}),n(\check{t})<0$.
Notice that
(A.22) $\Delta(t)=-m(t)n(t)+\frac{1-t}{t}\big{[}C_{b}m(t)-tC_{g}n(t)\big{]},$
and
(A.23) $\Delta^{\prime}(t)=\frac{1}{t}\big{[}C_{b}m(t)-C_{g}n(t)\big{]}.$
Recall that $\Delta^{\prime}(t^{\star})=0$. Therefore, by (A.23), $C_{b}m(t^{\star})=C_{g}n(t^{\star})$. Hence,
$m(t^{\star})=\frac{C_{b}m(t^{\star})-t^{\star}C_{g}n(t^{\star})}{(1-t^{\star})C_{b}}=\frac{t^{\star}}{(1-t^{\star})^{2}C_{b}}\big{[}\Delta(t^{\star})+m(t^{\star})n(t^{\star})\big{]}.$
Since $\Delta(t^{\star})>0$, and $m(t^{\star})n(t^{\star})=\frac{C_{b}}{C_{g}}(m(t^{\star}))^{2}\geq 0$, we conclude that $m(t^{\star}),n(t^{\star})>0$. By monotonicity, $m(\hat{t}),n(\hat{t})>0$.
We now prove that $m(\check{t}),n(\check{t})$ are negative. First, we show that $m(\check{t})n(\check{t})>0$. Suppose that on the contrary, $m(\check{t})n(\check{t})\leq 0$. Since $\Delta^{\prime}(\check{t})>0$, (A.23) implies that $m(\check{t})\geq 0>n(\check{t})$ or $m(\check{t})>0\geq n(\check{t})$. Therefore, $C_{b}m(\check{t})-\check{t}C_{g}n(\check{t})>0$. Since $\Delta(\check{t})=0$, (A.22) implies that $m(\check{t})n(\check{t})>0$, a contradiction. Now suppose that $m(\check{t}),n(\check{t})$ are both positive. As $t\downarrow 0$, $n(t)$ goes to $\infty$- ∞. Therefore, there exists a unique $t^{\prime}\in(0,\check{t})$ such that $n(t^{\prime})=0$. Since $\Delta^{\prime}$ is decreasing, $\Delta^{\prime}(t^{\prime})>\Delta^{\prime}(\check{t})>0$. Hence, by (A.23), $m(t^{\prime})>0$, and by (A.22), $\Delta(t^{\prime})>0=\Delta(\check{t})$. This is a contradiction to $\Delta(t)$ being strictly increasing in $t$ on $(0,t^{\star})$ and that $t^{\prime}<\check{t}<t^{\star}$. Hence, $m(\check{t}),n(\check{t})<0$. ∎
Corollary A.3.
Consider the symmetric information benchmark in Section 3. Assume that $\mathbf{F}(C_{g},C_{b},\mu_{s},\mu_{r})>0$ and $C_{g}>\hat{K}(C_{b})$ and let $\hat{\pi}=(p,q)\in\Pi^{\circ}$ be the equilibrium outcome. $\hat{\mathbf{p}}(q)<p<\check{\mathbf{p}}(q)$.
Proof.
By Proposition 3.1, $M(p,q)>0>N(p,q)$. The result then follows Lemma A.2. ∎
Lemma A.4.
Let $C_{g}>\hat{K}(C_{b})$. Fixing any $\mu\in(0,1)$, $f((\hat{\mathbf{p}}(q),q),\mu)$ is positive and increasing in $q$.
Proof.
By (A.21), $\frac{q}{\hat{\mathbf{p}}(q)}=q+\hat{t}(1-q)>\hat{t}$ for all $q\in(0,1)$. Hence, $\hat{\mathbf{p}}(q)>f((\hat{\mathbf{p}}(q),q),\mu)>f((\hat{\mathbf{p}}(q),\hat{t}\hat{\mathbf{p}}(q)),\mu)$ for all $q\in(0,1)$. Taking $q\downarrow 0$, we have $\lim_{q\downarrow 0}f((\hat{\mathbf{p}}(q),q),\mu)=0$. Therefore, it suffices to show that $f((\hat{\mathbf{p}}(q),q),\mu)$ is increasing in $q$. By (A.13) and (A.14),
$\frac{\partial}{\partial q}f((\hat{\mathbf{p}}(q),q),\mu)$ $=\hat{\mathbf{p}}^{\prime}(q)\left[\mu-\mu C_{g}\ln\frac{\hat{\mathbf{p}}(q)(1-q)}{(1-\hat{\mathbf{p}}(q))q}-(1-\mu)C_{b}\frac{\hat{\mathbf{p}}(q)-q}{\hat{\mathbf{p}}(q)(1-\hat{\mathbf{p}}(q))}\right]$
$\qquad\qquad\qquad+\left[(1-\mu)+\mu C_{g}\frac{\hat{\mathbf{p}}(q)-q}{q(1-q)}+(1-\mu)C_{b}\ln\frac{\hat{\mathbf{p}}(q)(1-q)}{(1-\hat{\mathbf{p}}(q))q}\right],$
and by (A.21),
$\hat{\mathbf{p}}^{\prime}(q)=\frac{\hat{\mathbf{p}}(q)(1-\hat{\mathbf{p}}(q))}{q(1-q)}.$
Therefore,
(A.24) $\frac{\partial}{\partial q}f((\hat{\mathbf{p}}(q),q),\mu)=J(\hat{\mathbf{p}}(q),q)\mu+K(\hat{\mathbf{p}}(q),q)(1-\mu),$
where
$J(p,q):=\frac{p(1-p)}{q(1-q)}\left(1+C_{g}\left[\frac{p-q}{p(1-p)}-\ln\frac{p(1-q)}{(1-p)q}\right]\right),$
$K(p,q):=1-C_{b}\left[\frac{p-q}{q(1-q)}-\ln\frac{p(1-q)}{(1-p)q}\right].$
Substituting $\hat{\mathbf{p}}(q)$ using (A.21) and using $\frac{q}{\hat{\mathbf{p}}(q)}>\hat{t}$, we have
$\frac{q(1-q)}{\hat{\mathbf{p}}(q)(1-\hat{\mathbf{p}}(q))}J(\hat{\mathbf{p}}(q),q)=1+C_{g}\left[\frac{q}{\hat{\mathbf{p}}(q)}\frac{1-\hat{t}}{\hat{t}}+\ln\hat{t}\right]>1+C_{g}(1-\hat{t}+\ln\hat{t})=m(\hat{t}),$
and
$K(\hat{\mathbf{p}}(q),q)=1-C_{b}\left[\frac{\hat{\mathbf{p}}(q)}{q}(1-\hat{t})+\ln\hat{t}\right]>1-C_{b}\left(\frac{1}{\hat{t}}-1+\ln\hat{t}\right)=n(\hat{t}).$
By Lemma A.2, $m(\hat{t}),n(\hat{t})$ are positive. Hence, $J(\hat{\mathbf{p}}(q),q),K(\hat{\mathbf{p}}(q),q)$ are positive for all $q\in(0,1)$. Therefore, by (A.24), $f((\hat{\mathbf{p}}(q),q),\mu)$ is increasing in $q$. ∎
A.6. Proof of Lemma 4.2
Proof.
To obtain a contradiction, suppose that there exists a D1 equilibrium where both sender types $i$ and $j$ choose $\pi=(p,q)$. Without loss of generality, let $i$ be the highest sender type that chooses $\pi$, so the receiver’s interim belief $\beta(\pi)<\mu_{i}$. For the rest of the proof, we assume that $\frac{\partial}{\partial\mu}MRS(\pi|\mu)<0$. A similar argument applies if $\frac{\partial}{\partial\mu}MRS(\pi|\mu)>0$.
Notice that $\pi$ must be persuasive at $\beta(\pi)<\mu_{i}$. Therefore, $\frac{q}{p}\leq\mathbf{Q}(\beta(\pi))<\mathbf{Q}(\mu_{i})$. By continuity, there is a neighborhood of $\pi$ wherein all experiments are persuasive at belief $\mu_{i}$.
For each sender type $\theta$, denote by $\mathfrak{q}_{\theta}$ her indifference curve through $\pi$. That is,
$f((p^{\prime},\mathfrak{q}_{\theta}(p^{\prime})),\mu_{\theta})=f(\pi,\mu_{\theta})$
for all $p^{\prime}$. $\mathfrak{q}_{\theta}$ is well defined on a small neighborhood of $p$, and it is the solution to the initial value problem
(A.25) $\mathfrak{q}_{\theta}^{\prime}(p^{\prime})=MRS((p^{\prime},\mathfrak{q}_{\theta}(p^{\prime}))|\mu_{\theta})~{}\text{and}~{}\mathfrak{q}_{\theta}(p)=q.$
By single-crossing, $\mathfrak{q}_{\theta}(p^{\prime})$ is strictly decreasing in $\theta$ for all $p^{\prime}>p$.
Hence, we can find $\tilde{\pi}=(\tilde{p},\tilde{q})$ such that $\tilde{\pi}$ is persuasive at belief $\mu_{i}$, $\tilde{p}>p$, and $\tilde{q}\in(\mathfrak{q}_{i}(\tilde{p}),\mathfrak{q}_{i-1}(\tilde{p}))$. By construction, $f(\tilde{\pi},\mu_{i})>f((\tilde{p},\mathfrak{q}_{i}(\tilde{p})),\mu_{i})=f(\pi,\mu_{i})=v^{\star}_{i}$, and $f(\tilde{\pi},\mu_{k})<f((\tilde{p},\mathfrak{q}_{k}(\tilde{p})),\mu_{k})=f(\pi,\mu_{k})=v^{\star}_{k}$ for all $k<i$. That is, $\tilde{\pi}$ is a profitable deviation for the sender type $i$ if $\beta(\tilde{\pi})\geq\mu_{i}$, so $\beta(\tilde{\pi})<\mu_{i}$. However, since $\tilde{\pi}$ is strictly equilibrium dominated for all sender types $k<i$, the D1 criterion requires that $\beta(\tilde{\pi})\geq\mu_{i}$, a contradiction. ∎
A.7. Proof of Proposition 4.3
We divide the proof into a sequence of lemmas. Lemma A.5 shows that (iii) is satisfied in all D1 pooling equilibria. Lemma A.6 shows that an experiment satisfying (iii) is a pooling equilibrium outcome if and only if (i) and (ii) are satisfied. Lemma A.7 finds $\hat{q}$ such that a pooling equilibrium outcome $(\hat{\mathbf{p}}(q),q)$ is selected by D1 if and only if (iv) is satisfied.
Lemma A.5.
Let $\pi=(p,q)\in\Pi^{\circ}$ be a D1 pooling equilibrium outcome. $p=\hat{\mathbf{p}}(q)$.
Proof.
By Lemma 4.2, either $p=\hat{\mathbf{p}}(q)$ or $p=\check{\mathbf{p}}(q)$. Suppose by way of contradiction that $p=\check{\mathbf{p}}(q)$. Denote by $\mathfrak{q}_{\theta}$ the indifference curve of type $\theta$ sender through $\pi$. $\mathfrak{q}_{\theta}$ is well defined on a small neighborhood of $p$ and solves the initial value problem (A.25). By Proposition 4.1, $\mathfrak{q}_{\theta}(p^{\prime})$ is strictly decreasing in $\theta$ for all $p^{\prime}>p$. Moreover, $\pi$ is persuasive at belief $\mu_{0}<\mu_{N}$. Hence, in a small neighborhood of $\pi$, all experiments are persuasive at belief $\mu_{N}$. We can choose therein $\tilde{\pi}=(\tilde{p},\tilde{q})$ such that $\tilde{p}>p$ and $\tilde{q}\in(\mathfrak{q}_{N}(\tilde{p}),\mathfrak{q}_{N-1}(\tilde{p}))$. By construction, $f(\tilde{\pi},\mu_{N})>f((\tilde{p},\mathfrak{q}_{N}(\tilde{p})),\mu_{N})=f(\pi,\mu_{N})=v^{\star}_{N}$, and $f(\tilde{\pi},\mu_{\theta})<f((\tilde{p},\mathfrak{q}_{\theta}(\tilde{p})),\mu_{\theta})=f(\pi,\mu_{\theta})=v^{\star}_{\theta}$ for all $\theta<N$. Hence, the D1 criterion requires $\beta(\tilde{\pi})=\mu_{N}$. But this makes $\tilde{\pi}$ a profitable deviation for the highest sender type, a contradiction. Therefore, $p=\hat{\mathbf{p}}(q)$. ∎
Lemma A.6.
An experiment $\pi=(\hat{\mathbf{p}}(q),q)$ is a pooling equilibrium outcome if and only if: (i) $\frac{q}{\hat{\mathbf{p}}(q)}\leq\mathbf{Q}(\mu_{0})$; (ii) $f(\pi,\mu_{1})\geq V(\mu_{1},\mu_{1})$.
Proof.
Necessity. On the equilibrium path, $\beta(\pi)=\mu_{0}$. If, contrary to (i), $\pi$ is unpersuasive at belief $\mu_{0}$, the sender’s payoff is $-c(\pi|\mu_{\theta})<0$. Then the uninformative experiment is a profitable deviation, a contradiction. Therefore, (i) is satisfied.
Consider a deviation $\tilde{\pi}$ that is persuasive at belief $\mu_{1}$. The lowest sender type’s payoff from deviating to $\tilde{\pi}$ is $f(\tilde{\pi},\mu_{1})$. By definition, $V(\mu_{1},\mu_{1})=\sup_{\tilde{\pi}\in\tilde{\Pi}_{1}}f(\tilde{\pi},\mu_{1})$, where $\tilde{\Pi}_{1}$ denotes the set of all experiments that are persuasive at belief $\mu_{1}$. Hence, the lowest sender type does not have a profitable deviation only if (ii) is satisfied.
Sufficiency. Let $\pi=(\hat{\mathbf{p}}(q),q)$ satisfy conditions (i) and (ii). We first show that, given any sender type $\theta>1$, $f(\pi,\mu_{\theta})\geq V(\mu_{\theta},\mu_{1})$. If $V(\mu_{\theta},\mu_{1})=0$, this follows from Lemma A.4. If $V(\mu_{\theta},\mu_{1})>0$, by Lemma A.4, there exists $\bar{q}_{\theta}\in(0,1)$ such that $f(\pi,\mu_{\theta})\geq V(\mu_{\theta},\mu_{1})$ if and only if $q\geq\bar{q}_{\theta}$, and by Lemma A.1, there exists a unique $\hat{p}_{\theta}>0$ such that $f((\hat{p}_{\theta},\mathbf{Q}(\mu_{1})\hat{p}_{\theta}),\mu_{\theta})=V(\mu_{\theta},\mu_{1})$. By Proposition 3.1, $V(\mu_{1},\mu_{1})>0$, and there exists a unique $\hat{p}_{1}>\hat{p}_{\theta}$ such that $f((\hat{p}_{1},\mathbf{Q}(\mu_{1})\hat{p}_{1}),\mu_{1})=V(\mu_{1},\mu_{1})$. Since $f((p,\mathbf{Q}(\mu_{1})p),\mu_{1})$ is increasing in $p$ on $[0,\hat{p}_{1}]$, $f((\hat{p}_{\theta},\mathbf{Q}(\mu_{1})\hat{p}_{\theta}),\mu_{1})>0$. Therefore, by Lemma A.4, there exists $\bar{q}_{1}\in(0,1)$ such that $f(\pi,\mu_{1})\geq f((\hat{p}_{\theta},\mathbf{Q}(\mu_{1})\hat{p}_{\theta}),\mu_{1})$ if and only if $q\geq\bar{q}_{1}$. By our assumption that (ii) is satisfied, $f(\pi,\mu_{1})\geq V(\mu_{1},\mu_{1})>f((\hat{p}_{\theta},\mathbf{Q}(\mu_{1})\hat{p}_{\theta}),\mu_{1})$, so $q>\bar{q}_{1}$. We are done once we show that $\bar{q}_{1}>\bar{q}_{\theta}$.
By way of contradiction, suppose that $\bar{q}_{1}\leq\bar{q}_{\theta}$. Denote by $\mathfrak{q}_{1}$ and $\mathfrak{q}_{\theta}$ type 1 and type $\theta$ senders’ indifference curves through $(\hat{p}_{\theta},\mathbf{Q}(\mu_{1})\hat{p}_{\theta})$, respectively. By Corollary A.3, $\hat{\mathbf{p}}(\mathbf{Q}(\mu_{1})\hat{p}_{\theta})<\hat{p}_{\theta}<\check{\mathbf{p}}(\mathbf{Q}(\mu_{1})\hat{p}_{\theta})$. Therefore, by Proposition 4.1, $\mathfrak{q}_{\theta}(p^{\prime})<\mathfrak{q}_{1}(p^{\prime})$ for all $p^{\prime}$ such that $\hat{\mathbf{p}}(\mathfrak{q}_{1}(p^{\prime}))\leq p^{\prime}<\hat{p}_{\theta}$. Specifically, at $p^{\prime}=\hat{\mathbf{p}}(\bar{q}_{\theta})$, we have $\bar{q}_{\theta}=\mathfrak{q}_{\theta}(\hat{\mathbf{p}}(\bar{q}_{\theta}))<\mathfrak{q}_{1}(\hat{\mathbf{p}}(\bar{q}_{\theta}))$. But since $\bar{q}_{1}\leq\bar{q}_{\theta}$, by Lemma A.4, $f((\hat{\mathbf{p}}(\bar{q}_{\theta}),\bar{q}_{\theta}),\mu_{1})\geq f((\hat{\mathbf{p}}(\bar{q}_{1}),\bar{q}_{1}),\mu_{1})=f((\hat{p}_{\theta},\mathbf{Q}(\mu_{1})\hat{p}_{\theta}),\mu_{1})$, so $\bar{q}_{\theta}\geq\mathfrak{q}_{1}(\hat{\mathbf{p}}(\bar{q}_{\theta}))$, a contradiction. Therefore, $\bar{q}_{1}>\bar{q}_{\theta}$.
We now show by construction that $\pi$ is a pooling equilibrium outcome. Let $\pi_{\theta}=\pi$ for all $\theta\in\Theta$, $\beta(\pi)=\mu_{0}$, and $\beta(\tilde{\pi})=\mu_{1}$ for all $\tilde{\pi}\neq\pi$. Moreover, let $\hat{\beta}$ and $\mathbf{a}$ be such that $\hat{\beta}(\pi^{\prime},s)=\mathbf{B}(\beta(\pi^{\prime}),\pi^{\prime},s)$ for all $\pi^{\prime}\in\Pi$ and $s\in\{g,b\}$, and $\mathbf{a}(\pi^{\prime},s)=1$ if and only if $\hat{\beta}(\pi^{\prime},s)\geq\bar{\beta}$. To check that ($(\{\pi_{\theta}\}_{\theta\in\Theta},\mathbf{a},\beta,\hat{\beta})$ is an equilibrium, we only need to show that any deviation $\tilde{\pi}\neq\pi$ is not profitable for the sender. If $\tilde{\pi}$ is persuasive at belief $\mu_{1}$, the sender’s payoff from deviating to $\tilde{\pi}$ is at most $V(\mu_{\theta},\mu_{1})\leq f(\pi,\mu_{\theta})$. If $\tilde{\pi}$ is unpersuasive at belief $\mu_{1}$, the sender’s deviation payoff is $-c(\tilde{\pi}|\mu_{\theta})\leq 0$. Therefore, the sender does not have a profitable deviation. ∎
Lemma A.7.
There exists $\hat{q}$ such that a pooling equilibrium outcome $(\hat{\mathbf{p}}(q),q)$ is a D1 pooling equilibrium outcome if and only if $q\geq\hat{q}$.
Proof.
To shorten notation, let $\bar{Q}=\mathbf{Q}(\mu_{N})$. If $\frac{q}{\hat{\mathbf{p}}(q)}>\bar{Q}$ for all $q\in(0,1)$, by Lemma A.6, there is no D1 pooling equilibrium outcome, so the result is trivial. For the rest of the proof, we assume that there exists some $q$ such that $\frac{q}{\hat{\mathbf{p}}(q)}\leq\bar{Q}$.
Consider the type $\theta$ sender’s indifference curve through some experiment $(\hat{\mathbf{p}}(q),q)$, where $q$ satisfies $\frac{q}{\hat{\mathbf{p}}(q)}\leq\bar{Q}$. This indifference curve intersects the straight line $\frac{q}{p}=\bar{Q}$ twice. Let $(p_{1},\bar{Q}p_{1})$ and $(p_{2},\bar{Q}p_{2})$ be the intersections, with $p_{1}\leq\hat{\mathbf{p}}(q)<p_{2}$. That is, $p_{1}$ and $p_{2}$ solve
(A.26) $f((p,\bar{Q}p),\mu_{\theta})=f((\hat{\mathbf{p}}(q),q),\mu_{\theta}).$
This defines $p_{1}$ and $p_{2}$ as functions of $\mu_{\theta}$ and $q$.
We first prove that a pooling equilibrium outcome $\pi=(\hat{\mathbf{p}}(q),q)$ is a D1 pooling equilibrium outcome if and only if $p_{2}(\mu_{\theta},q)$ is weakly decreasing in $\theta$. To show necessity, suppose that $p_{2}(\mu_{i},q)<p_{2}(\mu_{j},q)$ for some $i<j$. Let $\mathfrak{q}_{\theta}$ be the indifference curves of sender type $\theta$ through $\pi$. By Proposition 4.1, $\mathfrak{q}_{i}<\mathfrak{q}_{j}$ on a small neighborhood to the right of $\hat{\mathbf{p}}(q)$, but $\mathfrak{q}_{j}(p_{2}(\mu_{i},q))<\bar{Q}p_{2}(\mu_{i},q)=\mathfrak{q}_{i}(p_{2}(\mu_{i},q))$. Therefore, there exists some $\pi^{\prime}=(p^{\prime},q^{\prime})$ such that $p^{\prime}\in(\hat{\mathbf{p}}(q),p_{2}(\mu_{i},q))$, $\mathfrak{q}_{i}(p^{\prime})=\mathfrak{q}_{j}(p^{\prime})=q^{\prime}$, and $MRS(\pi^{\prime}|\mu_{i})>MRS(\pi^{\prime}|\mu_{j})$. Since $\frac{q^{\prime}}{p^{\prime}}<\bar{Q}$, all experiments are persuasive at belief $\mu_{N}$ in a small neighborhood of $\pi^{\prime}$. And by Proposition 4.1, $\mathfrak{q}_{\theta}(\tilde{p})$ is decreasing in $\theta$ for all $\tilde{p}>p^{\prime}$. Therefore, we can find a deviation $\tilde{\pi}=(\tilde{p},\tilde{q})$ such that $\tilde{\pi}$ is persuasive at belief $\mu_{N}$, $\tilde{p}>p^{\prime}$ and $\tilde{q}\in(\mathfrak{q}_{N}(\tilde{p}),\mathfrak{q}_{N-1}(\tilde{p}))$. By construction, $f(\tilde{\pi},\mu_{N})>f((\tilde{p},\mathfrak{q}_{N}(\tilde{p})),\mu_{N})=f(\pi,\mu_{N})$, and $f(\tilde{\pi},\mu_{\theta})<f((\tilde{p},\mathfrak{q}_{\theta}(\tilde{p})),\mu_{\theta})=f(\pi,\mu_{\theta})$ for all $\theta<N$. Hence, the D1 criterion requires $\beta(\tilde{\pi})=\mu_{N}$. But this makes $\tilde{\pi}$ a profitable deviation for the highest sender type, a contradiction. Therefore, $p_{2}(\mu_{\theta},q)$ is weakly decreasing in $\theta$.
Conversely, if $p_{2}(\mu_{\theta},q)$ is weakly decreasing in $\theta$, we show that if $f(\tilde{\pi},\mu_{i})>f(\pi,\mu_{i})$ for some $i$ and $\tilde{\pi}$ that is persuasive at belief $\mu_{N}$, $f(\tilde{\pi},\mu_{j})>f(\pi,\mu_{j})$ for all $j<i$. Therefore, $\pi$ is the outcome of a D1 pooling equilibrium where the receiver’s interim beliefs are such that $\beta(\tilde{\pi})=\mu_{1}$ for all $\tilde{\pi}\neq\pi$ that is persuasive at belief $\mu_{N}$. Notice that $f(\tilde{\pi},\mu_{i})>f(\pi,\mu_{i})$ if and only if $\tilde{p}\in(p_{1}(\mu_{i},q),p_{2}(\mu_{i},q))$ and $\tilde{q}\in(\mathfrak{q}_{i}(\tilde{p}),\bar{Q}\tilde{p}]$. Fix $j<i$. By Proposition 4.1, $\mathfrak{q}_{i}(\tilde{p})>\mathfrak{q}_{j}(\tilde{p})$ for all $\tilde{p}<\hat{\mathbf{p}}(q)$. Moreover, $\mathfrak{q}_{i}>\mathfrak{q}_{j}$ on a small neighborhood to the right of $\hat{\mathbf{p}}(q)$ and at $p_{2}(\mu_{i},q)$. Therefore, $\mathfrak{q}_{i}(\tilde{p})>\mathfrak{q}_{j}(\tilde{p})$ for all $\tilde{p}\in(\hat{\mathbf{p}}(q),p_{2}(\mu_{i},q))$. Otherwise, $\mathfrak{q}_{i}=\mathfrak{q}_{j}$ has two solutions $p^{\prime}_{1}<p^{\prime}_{2}$ on $(\hat{\mathbf{p}}(q),p_{2}(\mu_{i},q))$, and $\frac{\partial}{\partial\mu}MRS((p^{\prime}_{1},\mathfrak{q}_{i}(p^{\prime}_{1}))|\mu)<0<\frac{\partial}{\partial\mu}MRS((p^{\prime}_{2},\mathfrak{q}_{i}(p^{\prime}_{2}))|\mu)$, which is a contradiction to Proposition 4.1. Therefore, for all $\tilde{p}\in(p_{1}(\mu_{i},q),p_{2}(\mu_{i},q))$, $\mathfrak{q}_{j}(\tilde{p})\leq\mathfrak{q}_{i}(\tilde{p})$. Hence, for all $\tilde{\pi}$ that is persuasive at belief $\mu_{N}$, $f(\tilde{\pi},\mu_{i})>f(\pi,\mu_{i})$ implies $f(\tilde{\pi},\mu_{j})>f(\pi,\mu_{j})$.
We now show that $p_{2}(\mu_{\theta},q)$ is weakly decreasing in $\theta$ if and only if $q\geq\hat{q}$ for some $\hat{q}$. This is done by showing that the solution to the following system of equations of $p$ and $q$
(A.27) $p_{2}(\mu,q)=p~{}\text{for all}~{}\mu\in[0,1]$
is unique. [^13] Notice that $f(\pi,\mu)$ is linear in $\mu$ for all $\pi\in\Pi$. Therefore, (A.27) is equivalent to $p_{2}(\mu_{i},q)=p_{2}(\mu_{j},q)=p$ for some $i\neq j$. If (A.27) does not have a solution, either $p_{2}(\mu_{\theta},q)$ is striclty decreasing in $\theta$ for all $q$, or it is never weakly decreasing. In the former case, $\hat{q}=0$; in the latter case, we can without loss let $\hat{q}=1$. Hence, we proceed assuming that (A.27) admits a solution. Moreover, letting $(\hat{p},\hat{q})$ be the solution to (A.27), $p_{2}(\mu,q)$ is strictly increasing in $\mu$ if $q<\hat{q}$, and it is strictly decreasing in $\mu$ if $q>\hat{q}$.
Let $(\hat{p},\hat{q})$ be a solution to (A.27). It is sufficient to show that
(A.28) $\frac{\partial^{2}}{\partial\mu\partial q}p_{2}(\mu,q)\Big{|}_{q=\hat{q}}<0.$
Evaluating the left-hand side of (A.26) at $p=p_{2}:=p_{2}(\mu,q)$ and taking derivative with respect to $q$,
$\frac{\partial}{\partial q}f((p_{2},\bar{Q}p_{2}),\mu)=[M(p_{2},\bar{Q}p_{2})\mu+N(p_{2},\bar{Q}p_{2})(1-\mu)]\frac{\partial p_{2}}{\partial q}.$
The derivative of the right-hand side of (A.26) with respect to $q$ is given already by (A.24). Therefore,
(A.29) $\frac{\partial p_{2}}{\partial q}=\frac{J(\hat{\mathbf{p}}(q),q)\mu+K(\hat{\mathbf{p}}(q),q)(1-\mu)}{M(p_{2},\bar{Q}p_{2})\mu+N(p_{2},\bar{Q}p_{2})(1-\mu)}.$
Taking derivative of (A.29) with respect to $\mu$, the left-hand side of (A.28) has the same sign as
(A.30) $N(\hat{p},\bar{Q}\hat{p})J(\hat{\mathbf{p}}(\hat{q}),\hat{q})-M(\hat{p},\bar{Q}\hat{p})K(\hat{\mathbf{p}}(\hat{q}),\hat{q}).$
Hence, it is equivalent to show that (A.30) is negative.
By Proposition 4.1, at any experiment $(p,q)$ such that $\hat{\mathbf{p}}(q)<p<\check{\mathbf{p}}(q)$, the slope of the indifference curve of a higher type sender is greater than that of a lower type sender. Hence, if the indifference curves of two sender types intersect twice at $(\hat{q},\hat{\mathbf{p}}(\hat{q}))$ and at $(\hat{p},Q\hat{p})$, it must be the case that $\hat{p}>\check{\mathbf{p}}(Q\hat{p})$. By Lemma A.2, $M(\hat{p},Q\hat{p}),N(\hat{p},Q\hat{p})<0$. By Lemma A.4, $J(\hat{\mathbf{p}}(\hat{q}),\hat{q}),K(\hat{\mathbf{p}}(\hat{q}),\hat{q})>0$. Lastly, we observe that
$N(p,q)J(p,q)-M(p,q)K(p,q)=\Delta(p,q).$
Hence, $N(\hat{\mathbf{p}}(\hat{q}),\hat{q})J(\hat{\mathbf{p}}(\hat{q}),\hat{q})-M(\hat{\mathbf{p}}(\hat{q}),\hat{q})K(\hat{\mathbf{p}}(\hat{q}),\hat{q})=0$. Therefore, it is equivalent to show
$\frac{N(\hat{p},\bar{Q}\hat{p})}{M(\hat{p},\bar{Q}\hat{p})}>\frac{K(\hat{\mathbf{p}}(\hat{q}),\hat{q})}{J(\hat{\mathbf{p}}(\hat{q}),\hat{q})}=\frac{N(\hat{\mathbf{p}}(\hat{q}),\hat{q})}{M(\hat{\mathbf{p}}(\hat{q}),\hat{q})}.$
After a change of variables, we can rewrite $\frac{N(p,q)}{M(p,q)}$ as a function of $t$, as is defined in (A.15), and $Q=\frac{q}{p}$,
$\frac{N(p,q)}{M(p,q)}=Q\underbrace{\frac{\left(1-C_{b}\left[\frac{1}{t}-1+\ln t\right]\right)}{1+C_{g}\left(1-t+\ln t\right)}}_{=:g(t)}.$
Let $\bar{t}=\frac{\bar{Q}(1-\hat{p})}{1-\bar{Q}\hat{p}}$ and $\hat{Q}=\frac{\hat{q}}{\hat{\mathbf{p}}(\hat{q})}$. We need to show that $\bar{Q}g(\bar{t})>\hat{Q}g(\hat{t})$. Since $\hat{Q}<\bar{Q}$, and $g$ is positive, it is sufficient to show $g(\bar{t})>g(\hat{t})$. Since $\hat{t}>\bar{t}$, the proof is complete once we show that $g(t)$ is decreasing in $t$.
Observe that
(A.31) $g^{\prime}(t)=\frac{1-t}{t^{2}\left(1+C_{g}\left(1-t+\ln t\right)\right)^{2}}\big{[}C_{b}-C_{g}t+C_{g}C_{b}(2(1-t)+(1+t)\ln t)\big{]}.$
Taking derivative of the bracket in (A.31) with respect to $t$ yields
$C_{g}\left[C_{b}\left(\ln t+\frac{1}{t}-1\right)-1\right],$
which is strictly decreasing in $t$ on $(0,1)$ and has a unique zero $\left(1+\hat{x}(C_{b})\right)^{-1}$, where $\hat{x}(C_{b})$ is the unique solution to (A.5) on the positive real line. That is, the bracket in (A.31) obtains its maximum at $t=\left(1+\hat{x}(C_{b})\right)^{-1}$. Moreover, $02(1-t)+(1+t)\ln t<$ 0 for all $t\in(0,1)$. Hence, the bracket in (A.31) is strictly decreasing in $C_{g}$. Therefore, substituting $t=\left(1+\hat{x}(C_{b})\right)^{-1}$ and $C_{g}=\hat{K}(C_{b})$, we obtain a strict upper bound of the bracket in (A.31)
(A.32) $C_{b}-\frac{\hat{K}(C_{b})}{1+\hat{x}(C_{b})}+C_{b}\frac{\hat{K}(C_{b})}{1+\hat{x}(C_{b})}\left[2\hat{x}(C_{b})-\left(2+\hat{x}(C_{b})\right)\ln(1+\hat{x}(C_{b}))\right].$
Substituting $\hat{K}(C_{b})$ using (A.4) and $\ln(1+\hat{x}(C_{b}))$ using (A.5), (A.32) becomes zero. That is, $g^{\prime}(t)<0$. ∎
A.8. Proof of Proposition 4.4
Proof.
The receiver’ equilibrium payoff
(A.33) $U^{p}(q)=\hat{\mathbf{p}}(q)\mu_{0}-q(1-\mu_{0})\frac{\beta^{\star}}{1-\beta^{\star}}.$
Substituting (A.21) in (A.33) and taking derivative with respect to $q$,
(A.34) $\frac{d}{dq}U^{p}(q)=\frac{\mu_{0}\hat{t}}{[q+\hat{t}(1-q)]^{2}}-(1-\mu_{0})\frac{\beta^{\star}}{1-\beta^{\star}},$
which is decreasing in $q$. Hence, $U^{p}(q)$ is concave in $q$.
By Proposition 4.3, $\frac{\bar{q}}{\hat{\mathbf{p}}(\bar{q})}=\mathbf{Q}(\mu_{0})$. Hence, the D1 equilibrium outcome $\bar{\pi}:=(\hat{\mathbf{p}}(\bar{q}),\bar{q})$ is persuasive at belief $\mu_{0}$, and $V(\mu_{0},\mu_{0})\geq f(\bar{\pi},\mu_{0})>0$. By Lemma A.1, there is a unique persuasion equilibrium of the no-signal benchmark where the sender chooses $\pi^{ns}=(p^{ns},\mathbf{Q}(\mu_{0})p^{ns})$. By Corollary A.3, $p^{ns}>\hat{\mathbf{p}}(\mathbf{Q}(\mu_{0})p^{ns})$, so $\frac{\mathbf{Q}(\mu_{0})p^{ns}}{\hat{\mathbf{p}}(\mathbf{Q}(\mu_{0})p^{ns})}>\mathbf{Q}(\mu_{0})$. Hence, $\mathbf{Q}(\mu_{0})p^{ns}>\bar{q}$, since $\frac{q}{\hat{\mathbf{p}}(q)}$ is increasing in $q$. That is, $\pi^{ns}$ is Blackwell more informative than $\bar{\pi}$. Hence, the receiver is strictly better off in the no-signal benchmark. Formally,
$U^{p}(\bar{q})=\frac{\bar{q}}{\mathbf{Q}(\mu_{0})}\mu_{0}[1-\mathbf{Q}(\beta^{\star})]<p^{ns}\mu_{0}[1-\mathbf{Q}(\beta^{\star})]=U^{ns}.$
A.9. Proof of Proposition 4.5
The proof uses the following lemma, which is proved at the end.
Lemma A.8.
Let $C_{g}>\hat{K}(C_{b})$. There exists a unique $\pi\in\Pi^{\circ}$ such that $f(\pi,\mu_{\theta})=0$ for all $\theta$. Moreover, $\frac{\partial}{\partial\mu}MRS(\pi|\mu)<0$.
Let $\pi^{\star}=(p^{\star},q^{\star})$ be the unique $\pi$ in Lemma A.8, and $\bar{\mu}$ be such that $\mathbf{Q}(\bar{\mu})=\frac{q^{\star}}{p^{\star}}$. We show that a trivial D1 equilibrium exists if and only if conditions (i) and (ii) are satisfied.
Suppose that a trivial D1 equilibrium exists, but, contrary to the claim, (i) is not satisfied. Then there exists an experiment $\hat{\pi}_{1}$ that gives the lowest sender type a positive payoff $V(\mu_{1},\mu_{1})$ regardless of the receiver’s belief, hence it is a profitable deviation, a contradiction. Suppose that (ii) is not satisfied, i.e., $\frac{q^{\star}}{p^{\star}}<\mathbf{Q}(\mu_{N})$. Then in a neighborhood of $\pi^{\star}$, all experiments are persuasive at belief $\mu_{N}$. We can thus choose in there an experiment $\tilde{\pi}$ such that deviating to $\tilde{\pi}$ is profitable for the highest sender type but strictly equilibrium dominated for all other types of the sender. Hence, the trivial equilibrium does not satisfy the D1 criterion, a contradiction. Therefore, conditions (i) and (ii) must be satisfied.
Conversely, suppose that (i) and (ii) are satisfied. By construction, an experiment $\pi=(p,q)$ that is persuasive at belief $\mu_{N}$ gives some sender type $\theta$ a positive payoff, i.e., $f(\pi,\mu_{\theta})>0$, only if $p<p^{\star}$. And by Lemma A.8, $f(\pi,\mu_{\theta^{\prime}})>0$ for all $\theta^{\prime}<\theta$. That is, the lowest sender type is the most keen to deviate. By (i), if the receiver’s off-path belief is such that $\beta(\pi)=\mu_{1}$ for all $\pi$, the lowest sender type does not have a profitable deviation from choosing the uninformative experiment. Therefore, there exists a trivial D1 equilibrium where $\beta(\pi)=\mu_{1}$ for all $\pi$ that is not the uninformative experiment.
Proof of Lemma A.8.
Since $f(\pi,\mu)$ is linear in $\mu$, $f(\pi,\mu_{\theta})=0$ for all $\theta$ is equivalent to the following system of equations
(A.35) $p-C_{g}\left[p\ln\frac{p}{q}-(1-p)\ln\frac{1-q}{1-p}\right]=0,$
(A.36) $q+C_{b}\left[q\ln\frac{p}{q}-(1-q)\ln\frac{1-q}{1-p}\right]=0.$
The left-hand side of (A.35) is increasing in $q$. Given any $p\in(0,1)$, it goes to negative infinity as $q$ goes to zero and equals $p$ when $q=p$. Therefore, (A.35) defines a function $\sigma_{1}:p\mapsto q$ such that $0<\sigma_{1}(p)<p$ for all $p\in(0,1)$. Similarly, (A.36) defines a function $\sigma_{0}:p\mapsto q$ on $(0,1)$. Both $\sigma_{1}$ and $\sigma_{0}$ are continuous, hence their limits are well defined at 0 and 1. As $p$ goes to 0, $\lim_{p\downarrow 0}\sigma_{1}(p)=\lim_{p\downarrow 0}\sigma_{0}(p)=0$. As $p$ goes to 1,
(A.37) $\lim_{p\uparrow 1}\sigma_{1}(p)=\exp\left(-\frac{1}{C_{g}}\right)<1=\lim_{p\uparrow 1}\sigma_{0}(p).$
The first equality in (A.37) is by observing that the left-hand side of (A.35) converges to $1+C_{g}\ln(q)$ as $p$ goes to 1. The second equality in (A.37) is by observing that, the limit resides in $[0,1]$, but for all $q<1$, the left-hand side of (A.36) goes to negative infinity as $p$ goes to 1.
We now show that $\lim_{p\downarrow 0}\sigma_{1}^{\prime}(p)>\lim_{p\downarrow 0}\sigma_{0}^{\prime}(p)$. Therefore, for small positive $p$, $\sigma_{1}(p)>\sigma_{0}(p)$, so a solution exists for the system of equations (A.35), (A.36). Notice that, intuitively, $\sigma_{1}$ is the sender’s zero-payoff curve if her prior is 1. Therefore, $V(1,\mu_{r})>0$ if and only if $\mathbf{Q}(\mu_{r})>\lim_{p\downarrow 0}\sigma_{1}^{\prime}(p)$. By Lemma A.1, $Q=\lim_{p\downarrow 0}\sigma_{1}^{\prime}(p)$ is the solution to
(A.38) $01+C_{g}\left(\ln Q+1-Q\right)=$ 0 .
Notice that $\ln Q+1-Q$ is increasing in $Q$ on $(0,1]$. As $Q$ goes to 0, it goes to negative infinity, and it equals 0 when $Q=1$. Therefore, (A.38) has a unique solution which is increasing in $C_{g}$. By substituting $C_{g}=\hat{K}(C_{b})$ using (A.4) and solving (A.38), we obtain a strict lower bound
$\lim_{p\downarrow 0}\sigma_{1}^{\prime}(p)>\frac{1}{1+\hat{x}(C_{b})}.$
Similarly, $\sigma_{0}$ is the sender’s zero-payoff curve if her prior is 0. Therefore, $\lim_{p\downarrow 0}\sigma_{1}^{\prime}(p)$ solves $Q-C_{b}(Q\ln Q+1-Q)=0$. By (A.5),
$\lim_{p\downarrow 0}\sigma_{0}^{\prime}(p)=\frac{1}{1+\hat{x}(C_{b})}.$
Let $\pi=(p,q)$ be a solution to (A.35), (A.36). We wan to show that $\sigma_{1}^{\prime}(p)<\sigma_{0}^{\prime}(p)$. This simultaneously implies uniqueness and that $\frac{\partial}{\partial\mu}MRS(\pi|\mu)<0$, since $\sigma_{1}^{\prime}(p)=MRS(\pi|1)$ and $\sigma_{0}^{\prime}(p)=MRS(\pi|0)$. Taking derivative of the left-hand side of (A.35) with respect to $p$ and evaluating at $\pi$ yields
$1-C_{g}\left[\ln\frac{p}{q}+\ln\frac{1-q}{1-p}\right]=C_{g}\frac{1}{p}\ln\frac{1-q}{1-p}.$
The derivative with respect to $q$ is $\frac{p-q}{q(1-q)}$. Hence,
$\sigma_{1}^{\prime}(p)=\frac{q(1-q)}{p(p-q)}\ln\frac{1-q}{1-p}.$
Similarly,
$\sigma_{0}^{\prime}(p)=\frac{q(p-q)}{p(1-p)}\left(\ln\frac{1-q}{1-p}\right)^{-1}.$
Therefore, $\sigma_{1}^{\prime}(p)<\sigma_{0}^{\prime}(p)$ is equivalent to
(A.39) $\left(\ln\frac{1-q}{1-p}\right)^{2}<\frac{(p-q)^{2}}{(1-p)(1-q)}.$
Notice that both sides of (A.39) are decreasing in $q$ on $(0,p]$, and they equal 0 at $q=p$. Hence, it is sufficient to show that the right-hand decreases faster, that is,
$-2\ln\frac{1-q}{1-p}>-\frac{1-q}{1-p}+\frac{1-p}{1-q}.$
This is true since $\frac{1-q}{1-p}>1$ and $02\ln x-x+\frac{1}{x}<$ 0 for all $x>1$. ∎
A.10. Proof of Proposition 5.1
Proof.
We first show that there exists a unique collection of experiments satisfying conditions (i) to (iii). We then show that $\{\pi_{\theta}\}_{\theta\in\Theta}$ is a D1 equilibrium outcome if and only if these conditions are satisfied. Lastly, we compare the receiver’s equilibrium payoff with that in the public-signal benchmark.
Existence and uniqueness. Let $\theta\geq 2$ be such that $V(\mu_{\theta},\mu_{\theta})>0$, and fix $v^{\star}_{\theta-1}\in[0,V(\mu_{\theta-1},\mu_{\theta-1})]$ such that $v^{\star}_{\theta-1}>0$ if $V(\mu_{\theta-1},\mu_{\theta-1})>0$. Consider the following maximization problem with a relaxed constraint
(A.40) $\max_{\pi=(p,q)}f(\pi,\mu_{\theta})$
$s.t.~{}\frac{q}{p}\leq\mathbf{Q}(\mu_{\theta})~{}\text{and}~{}f(\pi,\mu_{\theta-1})\leq v_{\theta-1}^{\star}.$
We are to show that (A.40) has a unique solution $(p_{\theta},q_{\theta})$, and $\frac{q_{\theta}}{p_{\theta}}=\mathbf{Q}(\mu_{\theta})$. Therefore, (A.40) is equivalent to (5.1). Moreover, $v^{\star}_{\theta}$ is bounded from above by $V(\mu_{\theta},\mu_{\theta})$ and, as we show later in the proof, strictly positive. Hence, by induction on $\theta$, we can solve a unique collection of experiments $\{\pi_{\theta}\}_{\theta\in\Theta}$ satisfying conditions (i) to (iii).
If $v^{\star}_{\theta-1}>0$, there exists a unique $\tilde{\mu}\in(0,\mu_{\theta-1}]$ such that $V(\mu_{\theta-1},\tilde{\mu})=v^{\star}_{\theta-1}$. If $v^{\star}_{\theta-1}=0$, $V(\mu_{\theta-1},\mu_{\theta-1})=0$. Therefore, $\mathbf{F}(C_{g},C_{b},\mu_{\theta-1},\mu_{\theta-1})<0$, and by assumption, $\mathbf{F}(C_{g},C_{b},\mu_{\theta},\mu_{\theta})>0$, where $\mathbf{F}$ is defined in Lemma A.1. Hence, there exists some $\tilde{\mu}\in[\mu_{\theta-1},\mu_{\theta}]$ such that $\mathbf{F}(C_{g},C_{b},\mu_{\theta-1},\tilde{\mu})\leq 0$ and $\mathbf{F}(C_{g},C_{b},\mu_{\theta},\tilde{\mu})>0$, i.e., $V(\mu_{\theta},\tilde{\mu})>0$, and $V(\mu_{\theta-1},\tilde{\mu})=0$. In either case, there exists a unique experiment $\tilde{\pi}$ that is persuasive at belief $\tilde{\mu}$ such that $f(\tilde{\pi},\mu_{\theta})=V(\mu_{\theta},\tilde{\mu})$, and $f(\tilde{\pi},\mu_{\theta-1})<V(\mu_{\theta-1},\tilde{\mu})=v^{\star}_{\theta-1}$. Hence, (A.40) is equivalent to
(A.41) $\max_{\pi\in\hat{\Pi}}f(\pi,\mu_{\theta})$
where $\hat{\Pi}=\{\pi=(p,q):\frac{q}{p}\leq\mathbf{Q}(\mu_{\theta}),f(\pi,\mu_{\theta-1})\leq v^{\star}_{\theta-1},f(\pi,\mu_{\theta})\geq f(\tilde{\pi},\mu_{\theta})\}$ is nonempty and compact. Moreover, $\hat{\Pi}$ does not contain $(0,0)$, so $f(\cdot,\mu_{\theta})$ is continuous on $\hat{\Pi}$. Therefore, (A.41) admits a solution and its value is at least $f(\tilde{\pi},\mu_{\theta})>0$.
Let $\pi_{\theta}=(p_{\theta},q_{\theta})$ be a solution of (A.41). Uniqueness is shown once we show that $\frac{q_{\theta}}{p_{\theta}}=\mathbf{Q}(\mu_{\theta})$, since by single-crossing, the intersection of $\hat{\Pi}$ and $\frac{q}{p}=\mathbf{Q}(\mu_{\theta})$ is connected, and $f((p,\mathbf{Q}(\mu_{\theta})p),\mu_{\theta})$ is concave in $p$. Suppose by way of contradiction that $\frac{q_{\theta}}{p_{\theta}}<\mathbf{Q}(\mu_{\theta})$. Then in a neighborhood of $\pi_{\theta}$, all experiments are persuasive at belief $\mu_{\theta}$. By single-crossing, we can find in there an experiment $\pi^{\prime}$ such that $f(\pi^{\prime},\mu_{\theta})>f(\pi_{\theta},\mu_{\theta})\geq f(\tilde{\pi},\mu_{\theta})$ and $f(\pi^{\prime},\mu_{\theta-1})<f(\pi_{\theta},\mu_{\theta-1})\leq v^{\star}_{\theta-1}$. That is, $\pi^{\prime}\in\hat{\Pi}$ and $f(\pi^{\prime},\mu_{\theta})>f(\pi_{\theta},\mu_{\theta})$. This is a contradiction to $\pi_{\theta}$ being a maximizer of (A.41).
Necessity. Let $\{\pi_{\theta}\}_{\theta\in\Theta}$ be a D1 equilibrium outcome. Assume by way of contradiction that $V(\mu_{\theta},\mu_{\theta})=0$, but $\pi_{\theta}$ is not the uninformative experiment for some sender type $\theta$. By Lemma 4.2, $\pi_{\theta}$ is only chosen by the sender type $\theta$, so it is persuasive at belief $\beta(\pi_{\theta})=\mu_{\theta}$. But $V(\mu_{\theta},\mu_{\theta})=0$, so $f(\pi_{\theta},\mu_{\theta})<0$. Therefore, the uninformative experiment is a profitable deviation for the sender, a contradiction. Hence, (i) is satisfied. Notice that the value of maximization problem (A.41), which equals $v^{\star}_{\theta}$ due to the equivalence between (A.41) and (5.1), is at least $f(\tilde{\pi},\mu_{\theta})>0$. Therefore, (ii) is implied by (iii).
We now show that $v^{\star}_{1}=V(\mu_{1},\mu_{1})$. If $V(\mu_{1},\mu_{1})=0$, $v_{1}^{\star}=0$ by (i). If $V(\mu_{1},\mu_{1})>0$, by Proposition 4.1, there exists a unique experiment $\hat{\pi}_{1}$ that is persuasive at belief $\mu_{1}$ such that $f(\hat{\pi}_{1},\mu_{1})=V(\mu_{1},\mu_{1})$. That is, the lowest sender type gets $V(\mu_{1},\mu_{1})$ from choosing $\hat{\pi}_{1}$ regardless of the receiver’s belief, so her equilibrium payoff $v^{\star}_{1}\geq V(\mu_{1},\mu_{1})$. On the other hand, $\pi_{1}$ is persuasive at belief $\mu_{1}$, so $v^{\star}_{1}=f(\pi_{1},\mu_{1})\leq V(\mu_{1},\mu_{1})$. Therefore, $v^{\star}_{1}=V(\mu_{1},\mu_{1})$.
For the rest of the proof, we denote by $v^{\star}_{\theta}$ the unique payoff of the sender type $\theta$ pinned down by (i) through (iii), and $\pi^{\star}_{\theta}=(p^{\star}_{\theta-1},\mathbf{Q}(\mu_{\theta})p^{\star}_{\theta})$ the solution to (A.41) for all $\theta\geq 2$ such that $V(\mu_{\theta},\mu_{\theta})>0$.
In order to show that the rest of (iii) holds, we prove the following lemma: $f(\pi^{\star}_{\theta},\mu_{\theta^{\prime}})\leq v^{\star}_{\theta^{\prime}}$ for all $\theta\geq 2$ such that $V(\mu_{\theta},\mu_{\theta})>0$ and $\theta^{\prime}<\theta$. Consider two cases. First, if $v^{\star}_{\theta-1}=0$, then by (i) and (ii), $v^{\star}_{\theta^{\prime}}=0$ for all sender types $\theta^{\prime}<\theta$. By single-crossing, $f(\pi^{\star}_{\theta},\mu_{\theta-1})\leq 0$ implies that $f(\pi^{\star}_{\theta},\mu_{\theta^{\prime}})\leq 0$ for all $\theta^{\prime}<\theta$. Second, if $\pi^{\star}_{\theta-1}$ is not the uninformative experiment, by single-crossing, $p>p^{\star}_{\theta-1}$ for all $\pi=(p,\mathbf{Q}(\mu_{\theta})p)\in\hat{\Pi}$. Otherwise, the indifference curve of the type $\theta$ sender through $\pi$ is lower than the indifference curve of the type $\theta-1$ sender through $\pi^{\star}_{\theta-1}$ on $[p,1)$. By construction, $\tilde{\mu}$ is such the indifference curve of the type $\theta-1$ sender through $\pi^{\star}_{\theta-1}$ is tangent to the straight line $\frac{q}{p}=\mathbf{Q}(\tilde{\mu})$. Therefore, for some $\pi^{\prime}=(p^{\prime},q^{\prime})$ such that $f(\pi^{\prime},\mu_{\theta})=f(\pi,\mu_{\theta})$, $\frac{q^{\prime}}{p^{\prime}}<\mathbf{Q}(\tilde{\mu})$. Hence, $f(\pi,\mu_{\theta})=f(\pi^{\prime},\mu_{\theta})<V(\mu_{\theta},\tilde{\mu})=f(\tilde{\pi},\mu_{\theta})$. That is, $\pi\notin\hat{\Pi}$, a contradiction. Hence, $p^{\star}_{\theta}>p^{\star}_{\theta-1}$. By single-crossing, $f(\pi^{\star}_{\theta},\mu_{\theta-1})\leq f(\pi^{\star}_{\theta-1},\mu_{\theta-1})$ implies that $f(\pi^{\star}_{\theta},\mu_{\theta^{\prime}})\leq f(\pi^{\star}_{\theta-1},\mu_{\theta^{\prime}})$ for all $\theta^{\prime}<\theta$. By induction on $\theta$, we conclude that $f(\pi^{\star}_{\theta},\mu_{\theta^{\prime}})\leq f(\pi^{\star}_{\theta^{\prime}},\mu_{\theta^{\prime}})=v^{\star}_{\theta^{\prime}}$ for all $\theta^{\prime}<\theta$.
We now show that condition (iii) holds. Suppose that $\pi_{\theta}\neq\pi^{\star}_{\theta}$ for some $\theta\geq 2$ such that $V(\mu_{\theta},\mu_{\theta})>0$. That is, $f(\pi^{\star}_{\theta},\mu_{\theta})>f(\pi_{\theta},\mu_{\theta})$. Let $q^{\prime}<q^{\star}_{\theta}$ be such that $f(\pi^{\prime},\mu_{\theta})>f(\pi_{\theta},\mu_{\theta})$, where $\pi^{\prime}=(p^{\star}_{\theta},q^{\prime})$. $\pi^{\prime}$ is persuasive at belief $\mu_{\theta}$, so it is a profitable deviation for the type $\theta$ sender if $\beta(\pi^{\prime})\geq\mu_{\theta}$. Moreover, $f(\pi^{\prime},\mu_{\theta^{\prime}})<f(\pi^{\star}_{\theta},\mu_{\theta^{\prime}})\leq v^{\star}_{\theta^{\prime}}$ for all $\theta^{\prime}<\theta-1$. That is, $\pi^{\prime}$ is strictly equilibrium dominated for all sender types $\theta^{\prime}\leq\theta$. Hence, the D1 equilibrium requires that $\beta(\pi^{\prime})\geq\mu_{\theta}$, but given this off-path belief, $\pi^{\prime}$ is a profitable deviation for the type $\theta$ sender, a contradiction. Hence, $\pi_{\theta}=\pi^{\star}_{\theta}$ for all $\theta$.
Sufficiency. Let $\{\pi_{\theta}\}_{\theta\in\Theta}$ satisfy conditions (i) to (iii). We are to construct a D1 equilibrium with outcome $\{\pi_{\theta}\}_{\theta\in\Theta}$. Let $\beta(\pi_{\theta})$ satisfy Bayes’ rule for all $\theta\in\Theta$, i.e., $\beta(\pi_{\theta})=\mu_{\theta}$ if $\pi_{\theta}$ is not the uninformative experiment. For all off-the-equilibrium-path experiments $\pi^{\prime}\notin\{\pi_{\theta}\}_{\theta\in\Theta}$, let $\hat{\Theta}({\pi^{\prime}})=\{\theta\in\Theta:f(\pi^{\prime},\mu_{\theta})>v^{\star}_{\theta}\}$. That is, $\hat{\Theta}({\pi^{\prime}})$ is the set of sender types who may find $\pi^{\prime}$ to be a profitable deviation if $\pi^{\prime}$ is persuasive at belief $\beta(\pi^{\prime})$. Let $\beta(\pi^{\prime})=\mu_{1}$ if $\hat{\Theta}(\pi^{\prime})=\emptyset$ and $\beta(\pi^{\prime})=\min\hat{\Theta}(\pi^{\prime})$ if $\hat{\Theta}(\pi^{\prime})$ is nonempty. Moreover, let $\hat{\beta}$ and $\mathbf{a}$ be such that $\hat{\beta}(\pi^{\prime},s)=\mathbf{B}(\beta(\pi^{\prime}),\pi^{\prime},s)$ for all $\pi^{\prime}\in\Pi$ and $s\in\{g,b\}$, and $\mathbf{a}(\pi^{\prime},s)=1$ if and only if $\hat{\beta}(\pi^{\prime},s)\geq\bar{\beta}$.
Notice that $D^{0}_{\theta}(\pi^{\prime})=D_{\theta}(\pi^{\prime})=[\mathbf{Q}^{-1}(\frac{q^{\prime}}{p^{\prime}}),\mu_{N}]$ for all $\pi^{\prime}=(p^{\prime},q^{\prime})\in\Pi^{\circ}$ and all $\theta\in\hat{\Theta}(\pi^{\prime})$, and $D_{\theta}(\pi^{\prime})=\emptyset$ for all $\theta\notin\hat{\Theta}(\pi^{\prime})$. Hence, the D1 criterion is satisfied. Hence, to check that ($(\{\Pi_{\theta}\}_{\theta\in\Theta},\mathbf{a},\beta)$ is a D1 equilibrium, we only need to check that no sender type has a profitable deviation. If $\beta(\pi^{\prime})=\mu_{1}$, since all sender types’ equilibrium payoffs $v^{\star}_{\theta}\geq V(\mu_{\theta},\mu_{1})$, $\pi^{\prime}$ is not a profitable deviation. If $\beta(\pi^{\prime})=\mu_{\theta}$ for some $\theta\geq 2$, $\pi^{\prime}$ is not persuasive at belief $\mu_{\theta}$. Suppose by way of contradiction that $\pi^{\prime}$ is persuasive at belief $\mu_{\theta}$. By definition, $f(\pi^{\prime},\mu_{\theta})>f(\pi_{\theta},\mu_{\theta})$, $f(\pi^{\prime},\mu_{\theta-1})\leq v^{\star}_{\theta-1}$. This is a contradiction to that $\pi_{\theta}$ is the solution to the maximization problem (5.1). Therefore, $\pi^{\prime}$ is not persuasive at belief $\mu_{\theta}$. Hence, the payoff from deviating to $\pi^{\prime}$ is negative for all sender types, so there is no profitable deviation. The receiver’s payoff. The sender chooses the same experiment in the D1 separating equilibrium and in the equilibrium of the public information benchmark if $\theta=1$ of if $V(\mu_{\theta},\mu_{\theta})=0$. If $V(\mu_{\theta},\mu_{\theta})>0$ for some $\theta>1$, denote by $\pi^{\star}_{\theta}=(p^{\star}_{\theta},\mathbf{Q}(\mu_{\theta})p^{\star}_{\theta})$ and $\hat{\pi}_{\theta}=(\hat{p}_{\theta},\mathbf{Q}(\mu_{\theta})\hat{p}_{\theta})$ the experiments chosen by the type $\theta$ sender in the D1 equilibrium and that in the public information benchmark, respectively. If $\hat{\pi}_{\theta}\in\hat{\Pi}$, where $\hat{\Pi}$ is defined in (A.41), $\pi^{\star}_{\theta}=\hat{\pi}_{\theta}$. If $\hat{\pi}_{\theta}\notin\hat{\Pi}$, $f((p,\mathbf{Q}(\mu_{\theta})),\mu_{\theta})$ is decreasing in $p$ at $p^{\star}_{\theta}$. Hence, $p^{\star}_{\theta}>\hat{p}_{\theta}$, that is, $\pi^{\star}_{\theta}$ is Blackwell more informative than $\hat{\pi}_{\theta}$. Therefore, conditional on the sender’s private information, the receiver’s payoff is weakly higher in the D1 equilibrium than in the public information benchmark. Taking expectation over the sender’s private information, we have $U^{s}\geq U^{ps}$. ∎
A.11. Proof of Proposition 6.1
Proof.
For $\underline{n}\leq n\leq\bar{n}$, let $z_{n}$ be the expected cost of acquiring additional signals before the difference between the number of $g$’s and $b$’s reaches either threshold, conditional on the state being good and the current difference being $n$. $\{z_{n}\}_{\underline{n}\leq n\leq\bar{n}}$ satisfies the recurrence relation
(A.42) $z_{n}=\bar{c}_{g}+\alpha z_{n+1}+(1-\alpha)z_{n-1}$
for all $\underline{n}<n<\bar{n}$, where $\bar{c}_{g}=\alpha c_{g}+(1-\alpha)c_{b}$, and the boundary conditions $z_{\underline{n}}=z_{\bar{n}}=0$. Rewriting (A.42) as
$\alpha\left(z_{n+1}+\frac{n+1}{2\alpha-1}\bar{c}_{g}\right)-\left(z_{n}+\frac{n}{2\alpha-1}\bar{c}_{g}\right)+(1-\alpha)\left(z_{n-1}+\frac{n-1}{2\alpha-1}\bar{c}_{g}\right)=0,$
we can solve that
$z_{n}=C_{1}x^{n}+C_{2}-\frac{n}{2\alpha-1}\bar{c}_{g},$
where $x=\frac{1-\alpha}{\alpha}$, and
$C_{1}=-\frac{\bar{c}_{g}}{2\alpha-1}\frac{\bar{n}-\underline{n}}{x^{\underline{n}}-x^{\bar{n}}},\quad C_{2}=\frac{\bar{c}_{g}}{2\alpha-1}\frac{\bar{n}x^{\underline{n}}-\underline{n}x^{\bar{n}}}{x^{\underline{n}}-x^{\bar{n}}}$
are constants solved using the boundary conditions. Hence, the expected cost of implementing the threshold stopping rule $\tau$ conditional on the good state, $\mathbb{E}[c_{g}n_{g}(h_{\tau})+c_{b}n_{b}(h_{\tau})|\omega=G]$, equals
$z_{0}=\frac{\bar{c}_{g}}{2\alpha-1}\frac{\bar{n}(x^{\underline{n}}-1)-\underline{n}(x^{\bar{n}}-1)}{x^{\underline{n}}-x^{\bar{n}}}.$
Similarly, conditional on the bad state, the expected cost of implementing $\tau$,
$\mathbb{E}[c_{g}n_{g}(h_{\tau})+c_{b}n_{b}(h_{\tau})|\omega=B]=\frac{\bar{c}_{b}}{2\alpha-1}\frac{\bar{n}(x^{\bar{n}}-x^{\bar{n}+\underline{n}})-\underline{n}(x^{\underline{n}}-x^{\bar{n}+\underline{n}})}{x^{\underline{n}}-x^{\bar{n}}}.$
Hence, the expected cost of implementing the strategy $\tau$, $\mathbb{E}[c_{g}n_{g}(h_{\tau})+c_{b}n_{b}(h_{\tau})]$, is
(A.43) $\frac{1}{2\alpha-1}\left[\mu_{0}\frac{\bar{n}(x^{\underline{n}}-1)-\underline{n}(x^{\bar{n}}-1)}{x^{\underline{n}}-x^{\bar{n}}}\bar{c}_{g}+(1-\mu_{0})\frac{\bar{n}(x^{\bar{n}}-x^{\bar{n}+\underline{n}})-\underline{n}(x^{\underline{n}}-x^{\bar{n}+\underline{n}})}{x^{\underline{n}}-x^{\bar{n}}}\bar{c}_{b}\right].$
We are to verify that (A.43) equals $\mathbb{E}[H(\mu_{0})-H(\hat{\mu})]$. First, notice that
$\ln\left(\frac{1-\mu_{n}}{\mu_{n}}\right)=\ln\left(\frac{1-\mu_{0}}{\mu_{0}}\right)+n\ln x$
for $n=\bar{n},\underline{n}$. And
$\frac{\mu_{0}-\mu_{\underline{n}}}{\mu_{\bar{n}}-\mu_{\underline{n}}}\mu_{\bar{n}}=\frac{\mu_{0}\left(1+\frac{1-\mu_{0}}{\mu_{0}}x^{\underline{n}}\right)-1}{\left(1+\frac{1-\mu_{0}}{\mu_{0}}x^{\underline{n}}\right)-\left(1+\frac{1-\mu_{0}}{\mu_{0}}x^{\bar{n}}\right)}=\frac{\mu_{0}(x^{\underline{n}}-1)}{x^{\underline{n}}-x^{\bar{n}}},$
$\frac{\mu_{\bar{n}}-\mu_{0}}{\mu_{\bar{n}}-\mu_{\underline{n}}}\mu_{\underline{n}}=\frac{1-\mu_{0}\left(1+\frac{1-\mu_{0}}{\mu_{0}}x^{\bar{n}}\right)}{\left(1+\frac{1-\mu_{0}}{\mu_{0}}x^{\underline{n}}\right)-\left(1+\frac{1-\mu_{0}}{\mu_{0}}x^{\bar{n}}\right)}=\frac{\mu_{0}(1-x^{\bar{n}})}{x^{\underline{n}}-x^{\bar{n}}}.$
Hence,
$\mathbb{E}\left[\hat{\mu}\ln\left(\frac{1-\hat{\mu}}{\hat{\mu}}\right)\right]=\mu_{0}\ln\left(\frac{1-\mu_{0}}{\mu_{0}}\right)+\mu_{0}\frac{\bar{n}(x^{\underline{n}}-1)-\underline{n}(x^{\bar{n}}-1)}{x^{\underline{n}}-x^{\bar{n}}}\ln x.$
Similarly,
$\mathbb{E}\left[(1-\hat{\mu})\ln\left(\frac{\hat{\mu}}{1-\hat{\mu}}\right)\right]=(1-\mu_{0})\ln\left(\frac{\mu_{0}}{1-\mu_{0}}\right)+(1-\mu_{0})\frac{\bar{n}(x^{\bar{n}}-x^{\bar{n}+\underline{n}})-\underline{n}(x^{\underline{n}}-x^{\bar{n}+\underline{n}})}{x^{\underline{n}}-x^{\bar{n}}}\ln x.$
Therefore, (A.43) equals $\mathbb{E}[H(\mu_{0})-H(\hat{\mu})]$. ∎
A.12. Proof of Proposition 6.2
Proof.
Let $\pi=(p,q)\in\Pi^{\circ}$. In the game with Shannon Entropy cost,
$\frac{\partial f(\pi,\mu)}{\partial p}=\mu\left[1-C\left(\ln\frac{1-r}{r}-\ln\frac{1-p}{p}\right)\right],$
$\frac{\partial f(\pi,\mu)}{\partial q}=(1-\mu)\left[1+C\left(\ln\frac{1-q}{q}-\ln\frac{1-r}{r}\right)\right],$
where $r:=\mu p+(1-\mu)q$. The marginal rate of substitution is therefore
(A.44) $MRS(\pi|\mu)=-\frac{\mu}{1-\mu}\left[1-\frac{CL}{R}\right],$
where $L:=\ln\frac{1-q}{q}-\ln\frac{1-p}{p}>0$, and $R:=1+C\left(\ln\frac{1-q}{q}-\ln\frac{1-r}{r}\right)>1$.
Taking derivative of (A.44) with respect to $\mu$,
(A.45) $\frac{\partial}{\partial\mu}MRS(\pi|\mu)=\frac{MRS(\pi|\mu)}{\mu(1-\mu)}-\frac{\mu}{1-\mu}\frac{C^{2}L}{R^{2}}\frac{p-q}{r(1-r)}.$
Hence, if $MRS(\pi|\mu)\leq 0$, $\frac{\partial}{\partial\mu}MRS(\pi|\mu)<0$. Therefore, $MRS(\pi|\mu)$ as a function of $\mu$ has at most one zero on $(0,1)$. Moreover, notice that $\lim_{\mu\downarrow 0}MRS(\pi|\mu)=0$, and $\lim_{\mu\uparrow 1}MRS(\pi|\mu)=-\infty$. Hence, $MRS(\pi|\mu)$ has a unique zero on $(0,1)$ if and only if
$\lim_{\mu\downarrow 0}\frac{\partial}{\partial\mu}MRS(\pi|\mu)=C\left(\ln\frac{1-q}{q}-\ln\frac{1-p}{p}\right)-1>0.$
That is,
$p>\tilde{\mathbf{p}}(q):=\frac{q}{q+(1-q)e^{\frac{1}{C}}}.$
If $p\leq\tilde{\mathbf{p}}(q)$, $MRS(\pi|\mu)$ is negative and decreasing in $\mu$ for all $\mu\in(0,1)$.
We now show that if $p>\tilde{\mathbf{p}}(q)$, $MRS(\pi|\mu)$ is single-peaked. Hence, it first increases and then decreases in $\mu$. This concludes statement (i). To shorten notation, we will briefly denote by $M,M^{\prime},M^{\prime\prime}$ the marginal rate of substitution $MRS(\pi|\mu)$ and its first and second order derivatives with respect to $\mu$, respectively. Multiplying (A.45) with $1-\mu$ and taking derivative with respect to $\mu$,
$(1-\mu)M^{\prime\prime}-M^{\prime}$ $=\frac{M^{\prime}}{\mu}-\frac{M}{\mu^{2}}-\frac{C^{2}L}{R^{2}}\frac{p-q}{r^{2}(1-r)^{2}}(r^{2}-2rq+q)+2\frac{C^{2}L}{R^{3}}\frac{p-q}{r^{2}(1-r)^{2}}(r-q)$
(A.46) $<\frac{M^{\prime}}{\mu}-\frac{M}{\mu^{2}}-\frac{C^{2}L}{R^{2}}\frac{p-q}{r^{2}(1-r)^{2}}(r^{2}-2rq+3q-2r)$
Suppose that at some $\mu\in(0,1)$, $M^{\prime}=0$. By (A.45), $M>0$, and
(A.47) $\frac{C^{2}L}{R^{2}}\frac{p-q}{r(1-r)}=\frac{M}{\mu^{2}}.$
Substituting $M^{\prime}=0$ and (A.47) in (A.46),
$(1-\mu)M^{\prime\prime}<-\frac{M}{\mu^{2}}\frac{3q-r-2rq}{r(1-r)}.$
Notice that
(A.48) $3q-r-2rq=-2(1-\mu)q^{2}+(4-\mu-2\mu p)q+\mu p.$
The right-hand side of (A.48) is increasing in $q$ on $(0,1)$, and it equals $\mu p>0$ at $q=0$. Hence, $03q-r-2rq>$ 0, and $M^{\prime\prime}<0$ at $\mu$. That is, $MRS(\pi|\mu)$ obtains its global maximum at $\mu$, and it has no local minimum. Therefore, $MRS(\pi|\mu)$ is single-peaked. Fix any $q\in(0,1)$ and $i<j$. Recall that, if $p\leq\tilde{\mathbf{p}}(q)$, $MRS(\pi|\mu)$ is decreasing in $\mu$, and observe from (A.45) that $\lim_{p\uparrow 1}\frac{\partial}{\partial\mu}MRS(\pi|\mu)=\infty$ for all $\mu\in(0,1)$. Hence, there exists some $\tilde{p}>\tilde{\mathbf{p}}(q)$ such that $MRS((\tilde{p},q)|\mu_{i})=MRS((\tilde{p},q)|\mu_{j})$. We are to show that
$\frac{\partial}{\partial p}\big{(}MRS((p,q)|\mu_{i})-MRS((p,q)|\mu_{j})\big{)}\Big{|}_{p=\tilde{p}}<0.$
Letting $\tilde{\mathbf{p}}_{i,j}:q\mapsto\tilde{p}$, statement (ii) then follows.
Taking derivative of (A.44) with respect to $p$,
(A.49) $\frac{\partial}{\partial p}MRS(\pi|\mu)=\frac{\mu}{1-\mu}\frac{C}{R}\left(\frac{1}{p(1-p)}-\frac{CL}{R}\frac{\mu}{r(1-r)}\right).$
Substituting $R$ in (A.49) using (A.44), we have
$\frac{\partial}{\partial p}MRS(\pi|\mu)=\frac{1}{L}\left(M+\frac{\mu}{1-\mu}\right)\left[\frac{1}{p(1-p)}-\left(1+\frac{1-\mu}{\mu}M\right)\frac{\mu}{r(1-r)}\right],$
where we again use $M$ to denote $MRS(\pi|\mu)$. Notice that $L>0$ is independent of $\mu$, and $MRS((\hat{p},q)|\mu_{i})=MRS((\hat{p},q)|\mu_{j})$. Hence, it is sufficient to show that
(A.50) $G(\mu):=\left(m+\frac{\mu}{1-\mu}\right)\left[\frac{1}{p(1-p)}-\left(1+\frac{1-\mu}{\mu}m\right)\frac{\mu}{r(1-r)}\right]$
is increasing given any $m,p$, and $q$. Taking derivative of (A.50) with respect to $\mu$,
$G^{\prime}(\mu)$ $=\frac{r^{2}-2rp+p}{r^{2}(1-r)^{2}}m^{2}-2\frac{r^{2}-2rq+q}{r^{2}(1-r)^{2}}m$
$\qquad\qquad\qquad\qquad\qquad+\frac{1}{(1-\mu)^{2}}\left[\frac{1}{p(1-p)}-\frac{\mu}{r(1-r)}\right]-\frac{\mu}{1-\mu}\frac{r^{2}-2rq+q}{r^{2}(1-r)^{2}}$
$=\frac{(mp-q)^{2}}{pr^{2}}+\frac{(m(1-p)-(1-q))^{2}}{(1-p)(1-r)^{2}}>0.$
That is, $G(\mu)$ is increasing. ∎